### v1

# Generative Emergent Communication: Large Language Model is a Collective World Model

Abstract
This study proposes a unifying theoretical framework called generative emergent communication (generative EmCom) that bridges emergent communication, world models, and large language models (LLMs) through the lens of collective predictive coding (CPC). The proposed framework formalizes the emergence of language and symbol systems through decentralized Bayesian inference across multiple agents, extending beyond conventional discriminative model-based approaches to emergent communication. This study makes the following two key contributions: First, we propose generative EmCom as a novel framework for understanding emergent communication, demonstrating how communication emergence in multi-agent reinforcement learning (MARL) can be derived from control as inference while clarifying its relationship to conventional discriminative approaches. Second, we propose a mathematical formulation showing the interpretation of LLMs as collective world models that integrate multiple agents’ experiences through CPC. The framework provides a unified theoretical foundation for understanding how shared symbol systems emerge through collective predictive coding processes, bridging individual cognitive development and societal language evolution. Through mathematical formulations and discussion on prior works, we demonstrate how this framework explains fundamental aspects of language emergence and offers practical insights for understanding LLMs and developing sophisticated AI systems for improving human-AI interaction and multi-agent systems.
Index Terms:
emergent communication, large language model, world model, generative model, symbol emergence.
I Introduction
Langauge evolves and changes over time as a result of decentralized human communications [1, 2, 3, 4]. Sentences are generated to describe a wide range of phenomena, including external events, emotions, and intentions. In particular, the system of language is not static but dynamic [5, 6, 7, 8]. As Peirce, the founder of semiotics, suggested, symbols, including language, can be characterized by a triadic relationship of sign, object, and interpretant [9, 10, 11]. Here, sign corresponds to words, sentences and other signals. In particular, the correspondence between sign and object, which is signified by a sign (i.e., signifier), is determined by an interpretant. In other words, the meaning of a sign, that is, language, depends on culture and context, and so on.
The series of studies on emergent communication (EmCom), which is also referred to as emergent language (EmLang), and symbol emergence has attempted to explain the emergence of language and sharing of meanings of language among agents [2, 12, 13, 14, 15, 16]. However, computational models and general theories that provide a comprehensive and integrative understanding of symbol and language emergence [17, 8] and address the interdependency between the two aspects are lacking: first, world modeling by agents, which depends on their embodiment and environmental adaptation [18, 19]; and second, language emergence, which involves the evolution of language structure reflecting structural knowledge of the world that becomes embedded in language through distributional semantics[20, 21]. Although we are entering the era of large language models (LLMs) and generative artificial intelligence (AI), a need for a theoretical framework that can explain the dynamic and semantic aspects of language emergence in embodied cognitive developmental systems still remains [22, 8, 23, 16].
Recently, numerous studies have discussed how LLMs possess a model of the world. LLMs learn the distribution of word or linguistic token sequences and become intelligent to the extent that they can solve a wide range of tasks, including question answering, machine translation, and conversations [24, 25]. These capabilities are widely believed to be based on distributional semantics [20, 22]. However, the emergence of a latent structure that enables distributional semantics in our human language is not yet uncovered, though the mechanism through which neural networks find the hidden structure of language is gradually being revealed. Even before the invention of LLMs, word2vec, skip-gram, or continuous bag of words language models could internally form relative knowledge of concepts, for instance, "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [26, 27]. Two language models independently modeling different languages can perform unsupervised machine translation through structural alignment to a certain extent [28]. Recently, it has become clearer that LLMs have much knowledge of the world, including color similarity, physical properties of materials, and spatial knowledge [29, 30, 31, 32, 33]. This implies that the capability of LLMs is based on the mysterious nature of language. We believe that uncovering this nature of human language is the crucial mission of studies on symbol/language emergence and EmCom/EmLang.
In related fields, for instance, evolutionary robotics, cognitive and developmental robotics, artificial life, computational linguistics, and machine learning (ML), a wide range of studies on language evolution, EmCom, and symbol emergence have been conducted [34, 14, 2, 23, 17]. Most of these studies relied on specific types of language games, for instance, referential, signaling, and naming games. Since the late 2010s, empowered by the representation learning and language modeling capability of deep neural networks (DNNs), studies on EmCom have been boosted again [35, 36, 37]. However, most studies only treated the formation of communication protocols in specific game settings.
These approaches somehow failed to construct a general framework capturing symbol emergence from the viewpoint perspective of general principles of environmental adaptation, such as the free-energy principle (FEP), predictive coding (PC), and world modeling. Recently, world models have garnered significant attention as representation-learning models that incorporate action outputs and temporal dynamics of agent– environment interactions [38, 18, 19]. This aligns with broader theoretical frameworks such as PC and the FEP. PC posits that the brain constantly predicts sensory information and updates its internal models to enhance predictability  [39], whereas FEP provides a more generalized framework explaining the self-organization of biological systems through minimization of free energy [40], which is associated with the idea of the Bayesian brain proposed by [41]. Notably, FEP extends beyond individual cognition to explain the self-organization of cognitive and biological systems in detail [42, 43, 44], making it a promising foundation for understanding symbol emergence at both individual and collective levels.
To address these challenges, Taniguchi proposed the collective predictive coding (CPC) hypothesis [16]. The idea is based on a generative view of cognitive systems. The CPC assumes that not only individual agents but also groups of agents committing to symbolic communications can be modeled as generative models. Regarding individual agents, the FEP and active inference provide a generative view of a cognitive system. In other words, the CPC hypothesis extends the idea of the FEP and PC to a societal level [45]. The CPC hypothesis argues that symbol/language emergence can be modeled as decentralized Bayesian inference of a shared latent representation in a hierarchical manner.
This study aims to provide a theoretical view that bridges the idea of world models, that is, internal models for the sensorimotor observations and dynamics, and the emergence of language. To this end, we extend the idea of the CPC hypothesis and formalize the framework of generative emergent communication (generative EmCom) by introducing a generative view to the existing framework of EmCom. This allows us to develop a more integrative general theory of multi-agent systems performing EmCom and symbol emergence systems.
The main contributions of this paper are as follows:
  1. 1.
    Propose a new framework of EmCom known as generative EmCom, demonstrating that the emergence of communication in multi-agent reinforcement learning (MARL) can be derived from the perspective of control as inference (CaI) within the framework of generative EmCom, and clarify the relationship between conventional discriminative EmCom and the new generative EmCom.
  2. 2.
    Present the mathematical framework showing how LLMs become a collective world model in a specific sense through collective PC.
The remainder of this paper is organized as follows: Section 2 introduces the framework of generative EmCom, formalizing the collective PC hypothesis and its mathematical foundations. Section 3 demonstrates the formulation of language games as decentralized Bayesian inference, with a focus on the Metropolis-Hastings naming game (MHNG). Section 4 explores the application of generative EmCom to MARL, showing how it facilitates cooperative behavior. Section 5 discusses the relationship between the proposed generative framework and conventional approaches to EmCom. Section 6 examines LLMs through the lens of collective world models, providing new insights into their capabilities and limitations. Finally, Section 7 concludes the paper and discusses future directions for research in this area.
II Generative EmCom
Figure 1: (Left) PGM representation of CPC in symbol emergence, that is, emergent communication or language. The top-down generative process corresponds to language understanding and prediction of observations, which is downward causation in the symbol emergence system. The bottom-up inference process corresponds to perception, representation learning, and communication, that is, language game, which is upward causation in the symbol emergence system. (Right) Overview of the CPC in a symbol emergence system illustrating the bidirectional process of language understanding and generation, mediated by inference through language games and representation learning.
II-A Collective predictive coding
The CPC hypothesis was proposed to explain the emergence of symbol systems, particularly language [16, 46]. The CPC extends the idea of PC and FEP from the individual cognition level to a societal level. Taniguchi et al. were the first to formulate the CPC from the perspective of FEP and active inference [45]. This suggests that symbol emergence follows a process of minimizing free energy across the entire multi-agent system.
Although PC theory suggests that individual brains constantly predict sensory information and update their internal representations, including world models. CPC suggests that a group of agents, for instance, a human society, predict sensory information of all of the agents and update its external representations, that is, symbol systems.
A question is raised. How can we update the external representations, e.g., language, while our brains are disconnected physically? The CPC hypothesis suggests that a type of language game performs a decentralized Bayesian inference among the group (e.g., [47, 48]). In this framework, language games (such as naming games) can be interpreted as implementing decentralized Bayesian inference of shared representations. A representative example is MHNG explained in Section II-C. The CPC hypothesis argues that symbol systems emerge as a result of decentralized Bayesian inference performed collaboratively by multiple agents.
Although the encoding of sensory information through internal representations is ensured by the plasticity of neural systems, the plasticity of external representations is guaranteed by the flexibility of our symbol systems. The arbitrariness of symbol systems is a widely recognized characteristic of symbols in semiotics [11]. Peirce referred to the process by which subjects assign meaning to symbols according to culture and context as the semiosis. Although our brains are physically and electrically separated, they are informationally connected through communication using a flexible symbol system. Therefore, with appropriate communication and symbol system update algorithms, we can encode information into the symbol system as an external representation. In fact, the CPC hypothesis can consider that humans collectively perform this action in language emergence.
This implies that language collectively encodes information about the world as observed by numerous agents through their sensory-motor systems. The CPC hypothesis study [16] did not provide a clear and detailed explanation regarding this point while proposing a new perspective on why LLMs seem to possess knowledge about the real world. This is one of the main topics of this paper.
Essentially, CPC hypothesizes that human language is formed through a process of collective PC, where the symbol system emerges to maximize the predictability of multi-modal sensory-motor information obtained by members of a society, that is, minimize the collective free energy of a group of agents. This approach provides a unified framework for understanding symbol emergence, language evolution, and the nature of linguistic knowledge from the perspective of environmental adaptation and brain science.
II-B Generative EmCom
Numerous studies on EmCom/EmLang have been conducted. Major approaches are language games, MARL, and iterated learning models (ILMs) [12, 6, 49, 50, 51]. Recently, EmCom studies based on language games, particularly referential games, have become a dominant approach in the research community of EmCom. These studies utilize the representation learning capability of DNNs, give agents language encoders and decoders, and make them learn to generate and interpret a complex sequence of tokens to identify its meaning. In a typical setting of referential games, a speaker agent encodes a given object to a sequence of tokens using a neural network, for instance, long short-term memory, and the listener agent decodes it and identifies the object among some other items. This approach considers EmCom as an optimization of communication protocol. Generally, this is in line with Shannon’s communication model [52]. This approach focuses less on the synergy between the perceptual and communicative aspects of language. This approach can be regarded as a discriminative model-based approach to EmCom/EmLang.
Unlike traditional EmCom models, generative EmCom models consider language emergence as a generative model. In the simplest case, a probabilistic graphical model (PGM) of generative EmCom based on CPC hypothesis can be described mathematically as follows:
Generative model:
$p\left(\{x^{k}\}_{k},\{z^{k}\}_{k},m\right)=p(m)\prod_{k}p\left(x^{k}\mid z^{k}\right)p\left(z^{k}\mid m\right)$ (1)
Inference model:
$q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)=q\left(m\mid\{z^{k}\}_{k}\right)\prod_{k}q\left(z^{k}\mid x^{k}\right)$ (2)
where $p(\cdot)$ and $q(\cdot)$ are assumed to be original probability density functions and approximate probability functions, respectively, following the conventional notation of variational inference. Variables and functions used in this paper are listed in Table I. In this framework, $q\left(z^{k}\mid x^{k}\right)$ is interpreted as the process of representation learning performed by agent $k$, whereas $q\left(m\mid\{z^{k}\}_{k}\right)$ is hypothesized to be derived through collaborative language games among the agents. The overall phenomenon of symbol emergence within the group can be considered as a distributed effort to approximate $q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)$ without centralized control.
Mathematically, the CPC frames the process of symbol emergence in terms of social representation learning, analogous to the formation of internal representations by individual brains. As we introduced in Section 3, the representation learning and the inference of latent variables can be naturally extended to multi-agent reinforcement learning (RL) involving EmCom. Additionally, the formulation of CPC from the perspective of active inference can make us understand that this generative EmCom is not only for creating communication protocol but also for the environmental adaptation as a system of agents. Similarly, FEP and active inference are the general explanatory theory of the brain’s adaptation to the environment. In summary, the framework of generative EmCom based on the CPC hypothesis offers an explanation for language evolution to serve environmental adaptation at a collective level.
Generative EmCom offers an intriguing dual perspective, depending on whether one focuses on the collective or the individual. From a holistic perspective, the system appears to perform centralized representation learning of $w$ across all agents. Notably, $o^{k}$ can be easily extended to include multiple sensory modalities. From the perspective of the system overseer, or generative EmCom, multi-agent symbol emergence becomes analogous to multi-modal representation learning of external representations. Conversely, from the perspective of an individual agent, the system engages in multi-modal representation learning, influenced by and contributing to the inference of a shared latent variable $m$, which parameterizes the prior distribution common to all agents. The equivalence of these two interpretations is the cornerstone of the generative EmCom perspective. In essence, generative EmCom can be formulated as the representation learning of shared external representations, mediated through the representation learning of internal representations by each agent, based on multi-agent and multi-modal sensory-motor information.
TABLE I: Nomenclature and parameter details
$m_{t}$ Message (shared latent variable) communicated between agents at time $t$
$o^{k}_{t}$ Optimality variable for the $k$-th agent (1: optimal, 0: not optimal)
$z^{k}_{t}$ Internal representation (e.g., state in RL) of the $k$-th agent at time $t$
$a^{k}_{t}$ Action of the $k$-th agent at time step $t$
$x^{k}_{t}$ Observations (or sensory inputs) of the $k$-th agent at time $t$
$\theta^{k}$ Global parameters of the internal models of the $k$-th agent, for instance, neural networks
$\phi^{k}$ Parameters of language model of the $k$-th agent.
$r^{k}$ Reward function for the $k$-th agent
$p(\cdot)$ Original probability density function (i.e., a generative model)
$q(\cdot)$ Approximate probability function (i.e., an inference model)
II-C Language Game as Decentralized Bayesian Inference
The hypothetical argument that language game can perform the decentralized Bayesian inference has a computational basis though whether actual language communication can realize such decentralized Bayesian inference in our human society is an open question. The MHNG is an instance of this idea.
The MHNG comprises the following steps:
  1. 1.
    Perception: Speaker and listener agents (Sp and Li) observe the $d$-th object, obtain $x^{Sp}_{d}$, and $x^{Li}_{d}$ infers their internal representations $z_{d}^{Sp}$ and $z_{d}^{Li}$, respectively.
  2. 2.
    MH communication: Speaker mentions the name $m_{d}^{Sp}$ of the $d$-th object by sampling it from $P(m_{d}|z_{d}^{Sp},\phi^{Sp})$. The listener determines whether it accepts the naming with probability $\gamma={\rm min}\left(1,\frac{P(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Sp})}{P(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Li})}\right)$.
  3. 3.
    Learning: After MH communication was performed for every object, the listener updates its global parameters $\theta^{Li}$ and $\phi^{Li}$.
  4. 4.
    Turn-taking: The speaker and listener alternate their roles and go back to (1).
It has been demonstrated that the MHNG is equivalent to the Metropolis-Hastings algorithm for inferring latent variables in a probabilistic generative model. This model conditions the internal representations $z^{k}$ of multiple agents, acting as representation learning machines, on a common external representation $m$. Although the original study assumed two agents and a categorical message $m$, the core PGM of generative EmCom underlying this theory does not make these assumptions. Consequently, this fundamental idea can be extended in various ways.
As described, unlike referential games, MHNG assumes a joint attention performed by two agents. This assumption may seem strange from the game-theoretic approach to EmCom, such as referential games. However, developmental studies suggest that joint attention is fundamental to human infant language acquisition. Human infants develop joint attention capabilities before vocabulary explosion occurs, and joint attention serves as a crucial foundation for language acquisition [7]. Gergely et al. highlighted that when children incorporate parental instructions into their learning process, they presuppose that parents have the intention to teach them [53].
Moreover, interactions modeled by referential games—where agents identify objects and receive feedback—are rarely observed in natural language acquisition in cognitive development. Unlike game-theoretic approaches, MHNG, as an example of generative EmCom, assumes cooperative behavior in symbol emergence, which aligns with scholarly observations about human language acquisition [7]. Thus, although conventional (discriminative) EmCom provides rational models for both cooperative and competitive games, the assumptions of generative EmCom better align with empirical findings in human communication behavior.
Figure 2: (Top) Overview of the MHNG process. In the game, which two agents (A and B) engage in, the agents perceive a target object with joint attention and form internal representations. One agent (speaker) utters a sign and the other (listener) determines whether to accept it. Thereafter, they take turns. (Bottom) PGM representation, which is assumed in the MHNG. The MHNG is proved to be an inference process in the representation learning in a collective multi-agent system [47].
The original idea of MH-based EmCom was introduced by Hagiwara et al. [48] and later generalized and formalized by Taniguchi et al. [47], who clarified its connection to representation learning. The concept of MHNG has since been extended and validated in various ways. Furukawa et al. [54] demonstrated that similar symbol emergence can be observed even when the relationship between $m$ and $z$ is head-to-head instead of tail-to-tail. The extension to multimodal sensory information was achieved by Hagiwara et al. [55], who demonstrated that MHNG can lead to symbol emergence even when agents have different sensory modalities. Their work also showed that modality information possessed by other agents, but not by oneself, can facilitate object category formation. Similar multimodal extensions have been explored with variational autoencoder (VAE)-based representation learning by Hoang et al. [56], though their work suggests that the mechanism for integrating multimodal sensory information influences symbol emergence.
Although these MHNG studies involve two agents, a mathematical extension to N-agent conditions was developed by Inukai et al. [57], who introduced a recursive structure in communication while maintaining its characteristics as distributed Bayesian inference. They argued that random partner selection for MHNG can be considered as a one-sample and limited-length approximation of this approach.
Although several generative EmCom studies define the sign $m$ as a categorical variable, this is not a necessary assumption of the framework. A recent study by Hoang et al. [58] demonstrates that sharing compositional word sequences is possible within the MHNG framework, similar to numerous EmCom studies. From Peirce’s semiotic perspective, signs include both compositional discrete sequences and continuous signs such as voice pitch and facial expressions. A recent study by You et al. [59] demonstrated that such continuous signs can emerge within the generative EmCom framework, whereas Saito et al. [60] modeled the emergence of compositional signs from continuous time-series information as signals.
To verify the CPC hypothesis, examining whether human sign acceptance rates in joint attention naming scenarios match MHNG predictions was necessary. An experimental semiotic study by Okumura et al. [61] demonstrated that the sign acceptance rate in MHNG effectively predicts human behavior.
MHNG serves as a basic language game to represent symbol emergence by realizing distributed Bayesian inference and inferring latent variables of the generative model, corresponding to language. The critical point is that the inference of the posterior distribution is performed in a decentralized manner through language games. This suggests that being based on the MH method is not a necessary condition. The MHNG-based approach represents an initial step in modeling symbol emergence (or language emergence, EmCom) as decentralized Bayesian inference. Future studies should explore the possibility of constructing generative EmCom models by distributing various inference methods (e.g., [62]).
III EmCom for Multi-agent Cooperation
III-A Generative EmCom for Multi-agent Reinforcement Learning
Figure 3: Graphical model of generating cooperative actions for two agents.
Communication and language are often considered to emerge to facilitate multi-agent cooperation. In recent years, studies on MARL with communication channels have been progressing. Initial methods in multi-agent deep RL include DIAL [63] and CommNet [64]. These methods connect the networks of agents through messages, enabling the learning of necessary messages for cooperative behavior through backpropagation. Additionally, multi-agent deep deterministic policy gradient (MADDPG), an extension of the DDPG [65] for MARL, has been proposed [66]. In these methods, the agents use messages sent from all other agents to learn their policies. Methods involving weighting or attention mechanisms have been developed to limit communication to only necessary agents [67, 68, 69, 70]. To compute efficient messages, graph neural networks (GNNs) are used [71, 72, 73, 74, 75]. In these studies, multiple agents are connected through a network and messages are inferred by making them differentiable variables through backpropagation. In other words, error information computed from the internal states of others is directly transmitted to oneself, which is an unnatural modeling from the perspective of communication among independent individuals.
In contrast, generative EmCom allows us to formulate MARL with emergent communication in a Bayesian manner by incorporating the idea of CaI [76], which is a theory to formulate RL as a PGM [77, 78]. Figure 3 shows a graphical model of cooperative action generation between two agents, and the details of each stochastic variable are listed in Table I. The behavior of each agent is generated through a Markov decision process with a prior variable $m_{t}$. The state $z^{k}_{t}$ of an agent at time $t$ is determined according to state $z^{k}_{t-1}$, action $a^{k}_{t-1}$, and message $m_{t}$, which is the shared latent variable:
$z^{k}_{t}\sim p(z^{k}_{t}|m_{t},z^{k}_{t-1},a^{k}_{t-1}).$ (3)
where $k\in\{A,B\}$ denotes an index of agents. The agent can indirectly infer the state of others through the message $m_{t}$ in a probabilistic manner.
The optimality variable $o^{k}_{t}\in\{0,1\}$ represents the state optimality of both agents: $1$ indicates that the state and action pair are on the optimal trajectories, whereas 00 indicates it is not. Note that optimality is a type of probabilistic interpretation of reward functions as shown below. In this model, two types of optimality exist: one for each individual agent and the other for the group of agents. The probability $p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})$ of this optimality variable is computed using reward function $r^{k}(z^{k}_{t},a^{k}_{t})$ as follows:
$p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})\propto\exp(r(z^{k}_{t},a^{k}_{t})).$ (4)
For the reward of the group of agents $r^{AB}$, the following approximate reward models are learned by each agent:
$p(o^{AB}_{t}=1|z^{k}_{t},a^{k}_{t})\approx\propto\exp(r^{AB}(z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})).$ (5)
[…]
Following the theory of CaI, the optimal state sequence for both agents can be calculated by inferring state $z_{t}$ and message $m_{t}$ under the condition that the value of the optimality variables is always 1, as if the two-agent system acts as a single agent:
$z^{A}_{t},m_{t}\sim p(z^{A}_{t},m_{t}|z^{B}_{t},o^{A}_{1:T}={\mathbf{1}},o^{AB}_{1:T}={\mathbf{1}}).$ (6)
However, this equation has two problems: it includes the internal state $z^{B}_{t}$ of others, which cannot be observed in practice, and deriving this probability distribution analytically is difficult. We solve these problems by alternately inferring the following two variables:
[…]
$m_{1:T}\sim p(m_{1:T}|z^{A}_{1:T},z^{B}_{1:T},o^{AB}_{1:T}={\mathbf{1}}){\rm:communication.}$
Equation (7) describes state planning, which can be computed based on the CaI framework [76]. Equation (8) describes the inference of the message and can be formulated using the MHNG proposed by Taniguchi et al. [55, 47], which allows both agents to infer messages through communication without observing each other’s internal states. Note that MHNG enables two agents to perform sampling in (8) without simultaneous observations of $z^{A}_{t},z^{B}_{t}$.
Thus, MHNG can be used not only for multimodal object categorization and naming but also for action coordination among multi-agents using PGM to formulate MARL, that is, modeling EmCom for multi-agent cooperation.
III-B Generative EmCom on World models
The generative EmCom for MARL can be extended to involve representation learning in the same way as we discussed in Section 2.3. This extension implies the integration of the idea of world models into generative EmCom.
The concept of a world model represents an internal model within an agent that captures the dynamics of environmental states, their responses to the actions of the agent, and their relationships with sensory inputs [79, 18, 19]. The concept of world model has its origins in the early days of AI and robotics studies [80]. Initial studies on ML investigated techniques for agents to autonomously construct and adapt their world models [81, 82]. Currently, the term generally refers to predictive frameworks [83], predominantly implemented using DNN architectures.
The concept of world models is closely related to the idea of PCs. PC is the idea that the brain constantly predicts sensory information and updates its internal models to minimize prediction errors. In the context of cognitive robotics and AI, world models provide the structure for representing and reasoning about the environment, whereas PC offers a mechanism for learning and updating these models based on sensory experiences. The free energy principle and active inference further unify these concepts, suggesting that both perception and action can be considered as processes of minimizing prediction errors or free energy. A theoretical connection exists between them[19] .
Generally, a theory of world models can be based on partially observable Markov decision process (POMDP). In this framework, the state $z_{t}$ at time $t$ is not directly observable by the agent. Instead, the agent receives an observation $x_{t}$, which is assumed to be generated from a latent state $z_{t}$. The agent’s actions $a_{t}$ influence the transition of states according to a probability distribution $p(z_{t+1}|z_{t},a_{t})$. The observation model is given by $p(x_{t}|z_{t})$. The goal of the agent is to learn these probability distributions and use them to make predictions and inferences about the environment. This can be formalized as:
State transition: $\quad z_{t+1}\sim p(z_{t+1}|z_{t},a_{t})$ (9)
Observation: $\quad x_{t}\sim p(x_{t}|z_{t})$ (10)
Inference: $\quad z_{t}\sim q(z_{t}|x_{1:t},a_{1:t-1})$ (11)
where $q(z_{t}|x_{1:t},a_{1:t-1})$ represents the agent’s belief about the current state given the history of observations and actions. Learning these models enables the agent to construct a comprehensive world model that can be used for planning and decision-making in complex, partially observable environments.
Incorporating world models into MARL based on generative EmCom is natural and reasonable. By estimating $z_{t}$ using the inference model of a world model $q(z_{t}|x_{1:t},a_{1:t-1})$, an agent can estimate its state. Agents are expected to perform EmCom in the same manner as described in Section 3.1. Figure 4 shows a PGM of generative EmCom on POMDPs. This demonstrates the theoretical connection between EmCom or language and world models. When considered in this way, CPC is the theory explaining that language is something that integrates world models of populated agents. This will be discussed in more detail in Section 5.
Figure 4: Graphical model of generative EmCom involving world models.
Several studies linking EmCom and world models have been conducted. Recent studies on EmLang have focused on MARL approaches, aiming to autonomously develop languages to solve tasks [2]. Because each agent typically has access only to its own observations, inferring the underlying state of the environment becomes crucial. Consequently, researchers have explored enabling RL agents to develop EmLangs for inter-agent communication while simultaneously learning world models to infer environmental states. Notably, not all MARL communication studies focus on developing symbol-based languages; some treat communication as simple information exchange. This distinction separates “ EmLang” from “learning tasks with communication.” [49]
Alexander et al. proposed a message-conditioned world model that improves the quality of communication among agents, facilitating the emergence of meaningful messages and enhancing the interpretability of the model [84]. Similarly, Kenzo Lobos-Tsunekawa et al. extended the widely used Dreamer world model by introducing a global world model, used exclusively during training. By incorporating information from all agents, they achieved improved sample efficiency through shared imagination [85]. Toledo et al. also used Dreamer but introduced GNNs to facilitate communication; however, their work is classified under ”learning tasks with communication” [86] because it does not aim to develop a model of EmLang.
IV Relationship with Conventional EmCom
Recent years have witnessed an increasing presence of research papers focused on EmCom at ML conferences such as NeurIPS and ICLR. This trend underscores the renewed interest in EmCom. In the field of EmCom, Lewis’ signaling game [87] framework is often used as a model of communication while other various formulations have also been proposed [35, 36, 37]. The game players, or agents, are optimized to achieve some goals, resulting in the communication protocols, which are referred to as EmLang. In the EmCom field, the extent to which EmLangs resemble human languages is frequently questioned from the perspectives of evolutionary and/or computational linguistics. For instance, questions have been raised about whether EmLangs have compositionality [88, 89] or whether they follow the well-known statistical universal properties of natural languages [90, 91, 92, 93]. Additionally, cases exist where cognitive constraints were modeled to induce characteristics of natural languages [94, 95].
The signaling game is a simple communication model that involves only a sender $S_{\boldsymbol{\phi}}(m|x)$ and a receiver $R_{\boldsymbol{\theta}}(x|m)$ and only allows unidirectional communication from the sender to the receiver. At each play, the signaling game proceeds as follows:
  1. 1.
    Observation: Sender $S_{\boldsymbol{\phi}}$ obtains an observation $x$, that is, $x\sim P(x)$.
  2. 2.
    Signaling: Sender $S_{\boldsymbol{\phi}}$ generates a message $m$ from the observation $x$, that is, $m\sim S_{\boldsymbol{\phi}}(m|x)$.
  3. 3.
    Reconstruction: Receiver $R_{\boldsymbol{\theta}}$ attempts to reconstruct the original observation $x$ from the message $m$ via $R_{\boldsymbol{\theta}}(x|m)$.
Sender $S_{\boldsymbol{\phi}}$ and receiver $R_{\boldsymbol{\theta}}$ are optimized via a gradient-based method toward successful communication. Conventionally, the objective function of the signaling game (to be maximized) is defined as follows [96, 97]:
$\mathcal{J}_{\textup{MI}}(\boldsymbol{\phi},\boldsymbol{\theta})\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}[\log R_{\boldsymbol{\theta}}(x|m)].$ (12)
We refer to $\mathcal{J}_{\textup{MI}}$ as the MI-maximizing objective function because it is known to be a variational lower bound (up to constant) of the following mutual information between $X$ and $M$ [98, 99]:
$I_{\boldsymbol{\phi}}(X;M)\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}\left[\log\frac{S_{\boldsymbol{\phi}}(m|x)}{\mathbb{E}_{P(x^{\prime})}[S_{\boldsymbol{\phi}}(m|x^{\prime})]}\right].$ (13)
This implies that the conventional signaling game in the field of EmCom has been formulated as a problem of maximizing the mutual information between $X$ and $M$, where $X$ and $M$ denote random variables corresponding to the realizations $x$ and $m$, respectively. In contrast, a recent study [100] has proposed a reformulation of the signaling game as a problem of maximizing the evidence lower bound (ELBO):
[…]
We refer to $\mathcal{J}_{\textup{ELBO}}$ as the ELBO-maximizing objective function, contrasting it with the MI-maximizing objective $\mathcal{J}_{\textup{MI}}$. By adopting the ELBO-maximizing objective function, we can introduce concepts from computational psycholinguistics into signaling games. To observe this, let us transform the ELBO maximizing objective function as follows:
$\begin{split}\mathcal{J}_{\textup{ELBO}}(\boldsymbol{\phi},\boldsymbol{\theta})&\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}[\underbrace{\log R_{\theta}(x|m)}_{\textup{communication}}+\beta\underbrace{\log P_{\theta}(m)}_{\begin{subarray}{c}\textup{(negative)}\\ \textup{surprisal}\end{subarray}}]\\ &\phantom{\mathrel{:=}}-\beta\underbrace{\mathbb{E}_{P(x)}\mathcal{H}(S_{\phi}(M|x))}_{\begin{subarray}{c}\textup{entropy}\\ \textup{maximizer}\end{subarray}},\end{split}$ (15)
Here, a term known as surprisal appears, which is a concept commonly used in computational psycholinguistics [101, 102, 103]. Surprisal is assumed to represent the cognitive load experienced by a listener/reader (or the receiver in the signaling game) when processing a sentence. Therefore, the ELBO-maximizing signaling game naturally models the trade-off between information transmission and surprisal.
A similar discussion involves modeling the trade-off between information transmission and efficiency, where studies have used the (variational) information bottleneck (IB, VIB) framework to model communication [104, 105, 106]. In fact, VIB is proven to be a generalization of (beta-)VAE [107, 108], resulting in similar models. In addition, a contemporary work [109] shows that a variant framework known as a referential game can also be reformulated with an ELBO-like objective, analogous in structure to that of a conditional VAE [110, 111].
In relation to discussions in the field of evolutionary linguistics, some studies have also incorporated ILM into the EmCom framework [112], which is another important research theme. The ILM is a framework that models generational changes, where supervised learning is repeatedly performed from parent agents to child agents. In the context of VAE and VIB, however, little discussion exists on modeling generational changes, and this remains a future challenge when considering generative symbol emergence.
The formulation of the signaling game presented in this section, that is, generative model-based re-formulation of conventional EmCom, shares some fundamental connections with the generative EmCom and CPC frameworks discussed in other sections. All these formulations can be interpreted as representation learning with messages (signals, signs, or sentences) serving as latent variables in generative models when using ELBO-type (or VIB-type) objectives. However, notable differences exist: Although the CPC hypothesis does not specify a particular method for posterior distribution estimation, the signaling game explicitly adopts amortized variational inference. Furthermore, the signaling game assumes asymmetric roles between senders and receivers, which is not necessarily the case in the CPC frameworks. Additionally, the extension of ELBO-type formulation to populated settings, where multiple agents interact, remains an open challenge in the signaling game framework and requires further investigation [^1] An ELBO-type formulation based on the FEP for populated settings of generative EmCom is described in [45]. However, an inference procedure, that is, a language game, for enabling agents to let a symbol system emerge has not been described..
V Large Language Model as Collective World models
Figure 5: Three levels of the relationship between language, perceptual, and action information. Left: (A) Image captioning and generation tasks corresponding language to a still image. (B) Video captioning and generation tasks corresponding language to a video, that is, a sequence of visual stimuli. (C) Action-dependent video captioning and generation corresponding language to dynamic perceptual and action information, which corresponds to a world model and a policy conditioned by language. Right: probabilistic generative models corresponding to each of (A) – (C).
V-A Our argument
Based on the discussion in Sections 2 and 3, we explain the main argument of this paper, stating that LLMs are collective world models.
To clarify the link between this argument and existing studies relating language and perceptual information, Figure 5 shows three levels of tasks relating language to visual information (an example of multimodal sensory information) and action streams, along with PGMs for each. These represent (A) image captioning and generation tasks, corresponding language to a still image [113, 114, 115]; (B) video captioning and generation tasks, corresponding language to a video, that is, a sequence of visual stimuli [116, 117]; and (C) action-dependent video captioning and generation, corresponding language to dynamic perceptual and action information, respectively [^2] Note that we redefine the generative process of world models assuming that the policies of agents that generate $a_{t}$ from $z_{t}$ are also a part of the generative process of this world model.. In particular, from the perspective of representation learning, (A) and (B) correspond to representation learning of images and videos, whereas (C) corresponds to world models explained in Section III-B. Language-conditioned world models and robotics foundation models, known as vision-language-action models, studied in robotics and autonomous vehicles correspond to (C) [118, 119, 120, 121, 122].
Corresponding to the three levels of captioning and generation tasks, we can consider generative EmCom. Figure 5 describes PGMs for generative EmCom corresponding to the three levels, respectively. Notably, they are only an instance of the generative model shown in the CPC hypothesis (Figure 1) and generalization of studies based on the MHNG described in Sections 2 and 3.
Figure 6: PGMs for generative EmCom corresponding to three levels of complexity: (A) image-, (B) video-, and (C) action- and video-based tasks. These models represent instances of the CPC hypothesis and generalizations of MHNG studies.
Here, $\theta^{k}$ represents the global parameters of the $k$-th agent, for instance, the parameters of neural networks for representation learning. In (C), the global parameter $\bar{\theta}^{k}$ includes not only the parameters of the prediction model of the world (i.e., Equations 9 and 10 ) but also those of policies $p(a_{t}|z_{t})$. Let us consider (C) as a generative model corresponding to a single super-agent, virtually. The super-agent has multimodal sensory information combining the observations $\{x^{k}_{1:t}\}_{k}$ and actions $\{a^{k}_{1:t}\}_{k}$ of every agent. In this case, (C) can be considered as a type of hierarchical world model combining $K$ agents and integrating their experiences. We can refer to the world model owned by the super-agent as a collective world model, where $\{\phi,\{\theta^{k}\}_{k}\}$ are its parameters.
From this perspective, the corpus data that comprise sentences uttered by agents are considered as a sample from approximate posterior distributions $m_{j}^{[i]}\sim q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})$, where $m_{j}^{[i]}$ denotes the $i$-th sentence describing the $j$-th observation $\{x^{k}_{j,1:t}\}_{k}$ and action $\{a^{k}_{j,1:t}\}_{k}$.
Language modeling implies approximating the distribution of word sequences $p(m|\phi^{k})$ in a corpus, where $\phi^{k}$ denotes a parameter of the language model of the $k$-th agent. This implies that the language model approximates the posterior distribution $q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})\approx p(m|\theta^{LM})$.
The rationale behind distributional semantics as modeled by LLMs can be comprehended through the CPC framework. When language emerges through CPC, it functions as a latent space that develops through representation learning based on the collective experiences of all agents — in this case, humans. These experiences encompass not only external sensory information but also interception and behavioral decision-making. Within this framework, sentences can be mathematically considered as samples drawn from the posterior distribution over messages $m$, formally expressed as $q(m|\{x^{k},a^{k}\}_{k})$, where $x^{k}$ and $a^{k}$ represent the observations and actions of agent $k$, respectively. This formulation naturally leads to the structural reflection of the distribution of $\{x^{k},a^{k}\}_{k}$ in that of $m$, analogous to the working mechanism of representation learning in individual agents.
A key insight is that language serves as a shared external representation that effectively combines the world models of multiple human agents. When examining the graphical model depicted in Figure 6 (C) as a cognitive model of a single ”super-agent,” we can interpret it as a factorized hierarchical world model incorporating $K$ distinct modalities. Through this lens, an LLM that models the distribution $q(m|\{x^{k},a^{k}\}_{k})$ can be considered as a collective world model — one that integrates the diverse experiences and knowledge of multiple agents into a unified representational space. This provides a theoretical foundation for understanding why LLMs trained on large-scale human-generated text corpora appear to capture rich world knowledge.
The emergence of these phenomena can be comprehensively understood through the lens of the CPC hypothesis. The hypothesis suggests that language evolution occurs through a process of CPC, where the symbol system (language) emerges to maximize the predictability of multimodal sensory-motor information obtained by members of a society. This theoretical framework helps explain why LLMs, trained on the linguistic outputs of this process, inherit structural properties that reflect collective human knowledge and experience.
The CPC hypothesis thus provides a principled explanation for the acquisition of world knowledge by language models without direct sensorimotor experience by tapping into the accumulated wisdom encoded in human language through this collective process.
V-B Existing argument
LLMs learn to predict text tokens in an autoregressive manner, and unlike human cognition or prediction, they are not designed to interact with environments or achieve specific goals beyond word prediction. Nevertheless, some studies have suggested that LLMs may implicitly contain a world model that simulates our physical world.
Hao et al. proposed “reasoning via planning,” where an LLM is reused as a world model to perform logical reasoning, employing Monte Carlo tree search to reason in a more planned manner [123]. This method demonstrated higher performance compared to the existing chain-of-thought approach. Li et al. used a GPT model trained only on Othello game logs and confirmed that the model made predictions based on an internal world model by performing tests using explicit intervention operations that changed the activation of the middle layer of the model [124]. This suggests that LLM does not simply rely on statistical inference but constructs some degree of world model in systems such as games. In contrast, Wang et al. developed a benchmark involving text games to test whether transition models or policies can be derived from language models and revealed that LLMs are not reliable world simulators [125].
Although these studies investigate whether LLMs can be used as world models, studies examining whether LLMs inherently possess such models are underway. Gurnee et al. demonstrated that LLMs acquire spatiotemporal linear representations of the real world, specifically identifying individual spatial and temporal neurons [21]. The authors argue that this corresponds to the basic elements of a world model, although they did not show the acquisition of a causal model that includes actions. Andreas argues that during the process of predicting the next token, a language model infers the internal states, such as the beliefs, desires, and intentions of the speaker, from the context and reflects them in text generation or speech directed at other agents, suggesting the presence of a world model within the linguistic space [126]. Previous studies also indicated the possibility of representing belief states and intentions during text modeling. However, this pertains to agents within the linguistic space and does not explain connections with data derived from the physical world.
[…]
VI Conclusion
This study proposed a theoretical framework that unified EmCom, world models, and LLMs through the lens of CPC. We introduced the concept of generative EmCom as an alternative formulation of the conventional EmCom, which is based on a discriminative model-based language game such as referential games, and described their relationships. The concept, generative EmCom, is based on the CPC hypothesis, which demonstrates the modeling of language emergence as decentralized Bayesian inference of shared latent representations. We showed the application of this framework to MARL and provided a novel perspective on LLMs as collective world models that integrate the diverse experiences and knowledge of multiple agents into a unified representational space.
The theoretical framework proposed here offers several important insights. First, it provides a principled explanation for how language models can acquire world knowledge without direct sensorimotor experience, by tapping into the accumulated wisdom encoded in human language through CPC. Second, it bridges the gap between individual cognitive development and collective language evolution by showing how both can be comprehended through the lens of representation learning and free energy minimization. Third, it suggests that the success of LLMs may be fundamentally linked to their ability to capture and integrate the collective world models acquired by humans, i.e., through embodied sensorimotor interactions with the world.
However, several limitations of the current work should be addressed. First, although we provide a theoretical framework, direct empirical evidence for the collective world model hypothesis remains limited. The relationship between neural representations in language models and human conceptual structures needs further investigation. Second, the proposed framework primarily focuses on the emergence of linguistic structure and meaning but does not fully address the emergence of pragmatic aspects of language use. Third, the current formulation may not fully capture the dynamic, interactive nature of human language evolution, for instance, language evolution over generations and through the interaction of several agents in an open world.
From the perspective of world models, understanding the influence of collective world models on the environmental adaptation of agents is also important. According to the proposed theoretical framework, a language formed by multiple agents with the same embodiment learning in the same environment should serve as an appropriate prior distribution for individual world models. In other words, EmLang should accelerate both world model learning and following environmental adaptation based on RL and other methods. Obtaining constructive evidence for this relationship is also crucial.
Our argument and theoretical framework initiate several promising directions for future studies. Although we have established the basic connections between emergent communication, world models, and LLMs, significant work remains to validate and extend these ideas. Key priorities include developing experimental paradigms to test the CPC hypothesis, creating more sophisticated computational models and implementations of generative EmCom, and investigating how language formed by embodied agents sharing similar environments could serve as an effective prior distribution for world model learning. From a theoretical perspective, a deeper mathematical formalization of the interaction between individual and collective learning processes will be crucial. In addition, this framework has important practical implications for the development of more capable multi-agent systems, improved human–AI interaction, and embodied AIs that rapidly adapt to their physical worlds and communicate their knowledge using EmLang. By pursuing these directions, we can work toward a more comprehensive understanding of language emergence and its role in environmental adaptation while advancing both the theoretical foundations and practical applications of AI systems.

### v2

@@
 Abstract
-This study proposes a unifying theoretical framework called generative emergent communication (generative EmCom) that bridges emergent communication, world models, and large language models (LLMs) through the lens of collective predictive coding (CPC).
-The proposed framework formalizes the emergence of language and symbol systems through decentralized Bayesian inference across multiple agents, extending beyond conventional discriminative model-based approaches to emergent communication.
-This study makes the following two key contributions: First, we propose generative EmCom as a novel framework for understanding emergent communication, demonstrating how communication emergence in multi-agent reinforcement learning (MARL) can be derived from control as inference while clarifying its relationship to conventional discriminative approaches.
-Second, we propose a mathematical formulation showing the interpretation of LLMs as collective world models that integrate multiple agents’ experiences through CPC.
-The framework provides a unified theoretical foundation for understanding how shared symbol systems emerge through collective predictive coding processes, bridging individual cognitive development and societal language evolution.
-Through mathematical formulations and discussion on prior works, we demonstrate how this framework explains fundamental aspects of language emergence and offers practical insights for understanding LLMs and developing sophisticated AI systems for improving human-AI interaction and multi-agent systems.
-Index Terms:
-emergent communication, large language model, world model, generative model, symbol emergence.
-I Introduction
-Langauge evolves and changes over time as a result of decentralized human communications [1, 2, 3, 4].
+Large Language Models (LLMs) have demonstrated a remarkable ability to capture extensive world knowledge, yet how this is achieved without direct sensorimotor experience remains a fundamental puzzle.
+This study proposes a novel theoretical solution by introducing the Collective World Model hypothesis.
+We argue that an LLM does not learn a world model from scratch; instead, it learns a statistical approximation of a collective world model that is already implicitly encoded in human language through a society-wide process of embodied, interactive sense-making.
+To formalize this process, we introduce generative emergent communication (Generative EmCom), a framework built on the Collective Predictive Coding (CPC).
+This framework models the emergence of language as a process of decentralized Bayesian inference over the internal states of multiple agents.
+We argue that this process effectively creates an encoder-decoder structure at a societal scale: human society collectively encodes its grounded, internal representations into language, and an LLM subsequently decodes these symbols to reconstruct a latent space that mirrors the structure of the original collective representations.
+This perspective provides a principled, mathematical explanation for how LLMs acquire their capabilities.
+The main contributions of this paper are: 1) the formalization of the Generative EmCom framework, clarifying its connection to world models and multi-agent reinforcement learning, and 2) its application to interpret LLMs, explaining phenomena such as distributional semantics as a natural consequence of representation reconstruction.
+This work provides a unified theory that bridges individual cognitive development, collective language evolution, and the foundations of large-scale AI.
+keywords:
+emergent communication , large language model , world model , generative model , symbol emergence
+[…]
+[…]
+1 Introduction
+Large Language Models (LLMs) have achieved astonishing success, exhibiting a profound capacity for reasoning and knowledge retrieval across countless domains [12, 83].
+A central puzzle, however, lies in their apparent grasp of the structure of the physical world.
+LLMs, trained on vast corpora of text, are not designed to interact with an environment and lack any direct sensorimotor experience [4, 10].
+Yet, they seem to possess what could be described as a “world model.” How can an LLM acquire such a rich model of the world without ever having perceived or acted within it?
+This paper proposes a theoretical solution to this fundamental question.
+The debate over whether LLMs truly possess world models is active and ongoing.
+A growing body of evidence suggests that LLMs acquire surprisingly rich representations of the real world.
+For instance, some studies show that LLMs implicitly learn representations corresponding to vision and hearing just by reading text [133], and that specific spatiotemporal representations emerge within their activations [36].
+Furthermore, this internal knowledge proves functionally potent; LLMs can be used as world models for complex planning tasks [46], and their distributional semantics can be grounded in the physical world to guide robotic perception and action [93, 142].
+Conversely, other work has revealed that LLMs are not reliable world simulators, suggesting their internal models are brittle or fundamentally different from those of embodied agents [136].
+This conflicting evidence highlights a deep conceptual gap in our understanding.
+To clarify our argument, we distinguish between two concepts of a world model [119, 26]:
+  1. (1)
+    Type 1 World Model (Internal Model): A subjective, internal model that an agent learns through its own sensorimotor interactions with the environment to predict future states and plan actions.
+This corresponds to the agent’s Umwelt [135].
+  2. (2)
+    Type 2 World Model (Model of the World): An objective, structured representation of knowledge about the world, its entities, and their relations, which may not be tied to a single agent’s direct experience.
+We can frame this distinction as one between a subjective Type 1 World Model and an objective Type 2 World Model.
+Much of the confusion in current discourse arises from this distinction; while the term ”world model” in AI often refers to a Type 1 model, particularly in the context of model-based reinforcement learning and predictive coding [38, 40, 39, 35, 119], discussions about LLMs often imply they possess a Type 2 model.
+This presents a paradox: it is difficult to see how a disembodied LLM could form a subjective, embodied Type 1 model, yet it is equally unclear how it could acquire an objective Type 2 model without any access to the world.
+This paper proposes a solution to this paradox.
+We argue that the seemingly objective knowledge within an LLM is, in fact, a pseudo-objective structure encoded in the distributional semantics of language.
+This structure emerges as a result of aggregating countless subjective, Type 1 world models from a society of embodied agents.
+From the perspective of generative emergent communication (Generative EmCom) an LLM does not model the world directly, but rather models the collective of these Type 1 models as externalized in language.
+This leads to our central hypothesis:
+{mdframed}
+[linewidth=0.5pt, roundcorner=5pt, innertopmargin=5pt, innerbottommargin=5pt] The Collective World Model Hypothesis
+Human language is not merely a communication protocol but serves as an externalized representation of a collective world model, which emerges from the decentralized, interactive sense-making processes of an entire society of embodied agents.
+LLMs acquire their world knowledge by learning a statistical approximation of this collective world model encoded in text corpora.
+To provide a formal basis for this hypothesis, we introduce a new theoretical framework called Generative EmCom.
+This framework is built upon the Collective Predictive Coding (CPC) [117], which extends the principles of predictive coding, the free-energy principle (FEP) and the Bayesian brain hypothesis [50, 32, 34, 27] from individual brains to a societal level.
+Generative EmCom formalizes how a shared symbol system, i.e., language, emerges as multiple agents collectively seek to minimize their shared prediction errors about the world, a process we model as decentralized Bayesian inference.
+However, such an integrative theory has been largely missing.
+The series of studies on emergent communication (EmCom) and symbol emergence has attempted to explain the formation of language [98, 68].
+Yet, these approaches have often failed to bridge the gap between two interdependent aspects: first, the learning of an individual agent’s world model, which is grounded in its embodiment and environmental adaptation [35, 119]; and second, the collective emergence of a language that reflects this grounded knowledge.
+Our work aims to address this specific challenge by providing a unified framework.
+This study, therefore, makes two primary contributions.
+First, we formalize the Generative EmCom framework, clarifying its relationship with conventional approaches and demonstrating its utility in multi-agent systems.
+Second, using this framework, we provide a principled, mathematical interpretation of LLMs as collective world models, offering a coherent explanation for their otherwise mysterious capabilities.
+This unified perspective bridges the gap between EmCom, cognitive development, and the foundations of large-scale AI, opening new frontiers for research.
+Indeed, the theoretical framework proposed herein has already begun to inspire new concrete models for multi-agent coordination in dynamic environments [88] and reward-independent communication in multi-agent reinforcement learning (MARL) [141].
+The remainder of this paper is organized as follows: Section 2 reviews the current landscape of EmCom and world models, identifying the theoretical gap our work addresses.
+Section 3 details the theoretical framework of Generative EmCom.
+Section 4 explains how a collective world model can emerge from multi-agent interaction and cooperation.
+Section 5 applies this theory to reinterpret LLMs.
+Finally, Section 6 discusses the implications and limitations of our work, and Section 7 concludes the paper.
+2 The Landscape of EmCom and World Models
+Language evolves and changes over time as a result of decentralized human communications [24, 98, 112, 113].
 Sentences are generated to describe a wide range of phenomena, including external events, emotions, and intentions.
-In particular, the system of language is not static but dynamic [5, 6, 7, 8].
+In particular, the system of language is inherently dyanamics rather than static [137, 114, 126, 123].
 As Peirce, the founder of semiotics, suggested, symbols, including language, can be characterized by a triadic relationship of sign, object, and interpretant [95, 96, 18].
@@
 In other words, the meaning of a sign, that is, language, depends on culture and context, and so on.
-The series of studies on emergent communication (EmCom), which is also referred to as emergent language (EmLang), and symbol emergence has attempted to explain the emergence of language and sharing of meanings of language among agents [2, 12, 13, 14, 15, 16].
-However, computational models and general theories that provide a comprehensive and integrative understanding of symbol and language emergence [17, 8] and address the interdependency between the two aspects are lacking: first, world modeling by agents, which depends on their embodiment and environmental adaptation [18, 19]; and second, language emergence, which involves the evolution of language structure reflecting structural knowledge of the world that becomes embedded in language through distributional semantics[20, 21].
-Although we are entering the era of large language models (LLMs) and generative artificial intelligence (AI), a need for a theoretical framework that can explain the dynamic and semantic aspects of language emergence in embodied cognitive developmental systems still remains [22, 8, 23, 16].
-Recently, numerous studies have discussed how LLMs possess a model of the world.
-LLMs learn the distribution of word or linguistic token sequences and become intelligent to the extent that they can solve a wide range of tasks, including question answering, machine translation, and conversations [24, 25].
-These capabilities are widely believed to be based on distributional semantics [20, 22].
-However, the emergence of a latent structure that enables distributional semantics in our human language is not yet uncovered, though the mechanism through which neural networks find the hidden structure of language is gradually being revealed.
-Even before the invention of LLMs, word2vec, skip-gram, or continuous bag of words language models could internally form relative knowledge of concepts, for instance, "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [26, 27].
-Two language models independently modeling different languages can perform unsupervised machine translation through structural alignment to a certain extent [28].
-Recently, it has become clearer that LLMs have much knowledge of the world, including color similarity, physical properties of materials, and spatial knowledge [29, 30, 31, 32, 33].
-This implies that the capability of LLMs is based on the mysterious nature of language.
-We believe that uncovering this nature of human language is the crucial mission of studies on symbol/language emergence and EmCom/EmLang.
-In related fields, for instance, evolutionary robotics, cognitive and developmental robotics, artificial life, computational linguistics, and machine learning (ML), a wide range of studies on language evolution, EmCom, and symbol emergence have been conducted [34, 14, 2, 23, 17].
-Most of these studies relied on specific types of language games, for instance, referential, signaling, and naming games.
-Since the late 2010s, empowered by the representation learning and language modeling capability of deep neural networks (DNNs), studies on EmCom have been boosted again [35, 36, 37].
-However, most studies only treated the formation of communication protocols in specific game settings.
-These approaches somehow failed to construct a general framework capturing symbol emergence from the viewpoint perspective of general principles of environmental adaptation, such as the free-energy principle (FEP), predictive coding (PC), and world modeling.
-Recently, world models have garnered significant attention as representation-learning models that incorporate action outputs and temporal dynamics of agent– environment interactions [38, 18, 19].
-This aligns with broader theoretical frameworks such as PC and the FEP.
-PC posits that the brain constantly predicts sensory information and updates its internal models to enhance predictability  [39], whereas FEP provides a more generalized framework explaining the self-organization of biological systems through minimization of free energy [40], which is associated with the idea of the Bayesian brain proposed by [41].
-Notably, FEP extends beyond individual cognition to explain the self-organization of cognitive and biological systems in detail [42, 43, 44], making it a promising foundation for understanding symbol emergence at both individual and collective levels.
-To address these challenges, Taniguchi proposed the collective predictive coding (CPC) hypothesis [16].
-The idea is based on a generative view of cognitive systems.
-The CPC assumes that not only individual agents but also groups of agents committing to symbolic communications can be modeled as generative models.
-Regarding individual agents, the FEP and active inference provide a generative view of a cognitive system.
-In other words, the CPC hypothesis extends the idea of the FEP and PC to a societal level [45].
-The CPC hypothesis argues that symbol/language emergence can be modeled as decentralized Bayesian inference of a shared latent representation in a hierarchical manner.
-This study aims to provide a theoretical view that bridges the idea of world models, that is, internal models for the sensorimotor observations and dynamics, and the emergence of language.
-To this end, we extend the idea of the CPC hypothesis and formalize the framework of generative emergent communication (generative EmCom) by introducing a generative view to the existing framework of EmCom.
-This allows us to develop a more integrative general theory of multi-agent systems performing EmCom and symbol emergence systems.
-The main contributions of this paper are as follows:
+2.1 Conventional Approaches to EmCom and Their Limits
+The study of how communication protocols and symbolic language emerge from multi-agent interaction, a field known as EmCom, has been explored through several major approaches, including language games, MARL, and iterated learning models (ILMs) [97, 68, 114, 146, 11].
+A central theme in this research, particularly in studies employing language games and powered by deep neural networks [31, 84, 55], is the extent to which the resulting protocols, or emergent languages (EmLangs), resemble human language.
+For instance, questions have been raised about whether EmLangs exhibit compositionality [65, 14], follow well-known statistical properties of natural languages [15, 105, 131, 129], or can be shaped by cognitive constraints [104, 57].
+A foundational paradigm for many of these investigations is the Lewis signaling game [73].
+This approach, often aligned with Shannon’s information theory [110], can be regarded as a discriminative model for optimizing a communication protocol.
+The signaling game is a simple communication model that involves only a sender $S_{\boldsymbol{\phi}}(m|x)$ and a receiver $R_{\boldsymbol{\theta}}(x|m)$ and only allows unidirectional communication from the sender to the receiver.
+At each play, the signaling game proceeds as follows:
   1. 1.
-    Propose a new framework of EmCom known as generative EmCom, demonstrating that the emergence of communication in multi-agent reinforcement learning (MARL) can be derived from the perspective of control as inference (CaI) within the framework of generative EmCom, and clarify the relationship between conventional discriminative EmCom and the new generative EmCom.
+    Observation: Sender $S_{\boldsymbol{\phi}}$ obtains an observation $x$, that is, $x\sim p(x)$.
   2. 2.
-    Present the mathematical framework showing how LLMs become a collective world model in a specific sense through collective PC.
-The remainder of this paper is organized as follows: Section 2 introduces the framework of generative EmCom, formalizing the collective PC hypothesis and its mathematical foundations.
-Section 3 demonstrates the formulation of language games as decentralized Bayesian inference, with a focus on the Metropolis-Hastings naming game (MHNG).
-Section 4 explores the application of generative EmCom to MARL, showing how it facilitates cooperative behavior.
-Section 5 discusses the relationship between the proposed generative framework and conventional approaches to EmCom.
-Section 6 examines LLMs through the lens of collective world models, providing new insights into their capabilities and limitations.
-Finally, Section 7 concludes the paper and discusses future directions for research in this area.
-II Generative EmCom
-Figure 1: (Left) PGM representation of CPC in symbol emergence, that is, emergent communication or language.
-The top-down generative process corresponds to language understanding and prediction of observations, which is downward causation in the symbol emergence system.
-The bottom-up inference process corresponds to perception, representation learning, and communication, that is, language game, which is upward causation in the symbol emergence system.
-(Right) Overview of the CPC in a symbol emergence system illustrating the bidirectional process of language understanding and generation, mediated by inference through language games and representation learning.
-II-A Collective predictive coding
-The CPC hypothesis was proposed to explain the emergence of symbol systems, particularly language [16, 46].
-The CPC extends the idea of PC and FEP from the individual cognition level to a societal level.
-Taniguchi et al.
-were the first to formulate the CPC from the perspective of FEP and active inference [45].
-This suggests that symbol emergence follows a process of minimizing free energy across the entire multi-agent system.
-Although PC theory suggests that individual brains constantly predict sensory information and update their internal representations, including world models.
-CPC suggests that a group of agents, for instance, a human society, predict sensory information of all of the agents and update its external representations, that is, symbol systems.
-A question is raised.
-How can we update the external representations, e.g., language, while our brains are disconnected physically?
-The CPC hypothesis suggests that a type of language game performs a decentralized Bayesian inference among the group (e.g., [47, 48]).
-In this framework, language games (such as naming games) can be interpreted as implementing decentralized Bayesian inference of shared representations.
-A representative example is MHNG explained in Section II-C.
-The CPC hypothesis argues that symbol systems emerge as a result of decentralized Bayesian inference performed collaboratively by multiple agents.
-Although the encoding of sensory information through internal representations is ensured by the plasticity of neural systems, the plasticity of external representations is guaranteed by the flexibility of our symbol systems.
-The arbitrariness of symbol systems is a widely recognized characteristic of symbols in semiotics [11].
-Peirce referred to the process by which subjects assign meaning to symbols according to culture and context as the semiosis.
-Although our brains are physically and electrically separated, they are informationally connected through communication using a flexible symbol system.
-Therefore, with appropriate communication and symbol system update algorithms, we can encode information into the symbol system as an external representation.
-In fact, the CPC hypothesis can consider that humans collectively perform this action in language emergence.
-This implies that language collectively encodes information about the world as observed by numerous agents through their sensory-motor systems.
-The CPC hypothesis study [16] did not provide a clear and detailed explanation regarding this point while proposing a new perspective on why LLMs seem to possess knowledge about the real world.
-This is one of the main topics of this paper.
-Essentially, CPC hypothesizes that human language is formed through a process of collective PC, where the symbol system emerges to maximize the predictability of multi-modal sensory-motor information obtained by members of a society, that is, minimize the collective free energy of a group of agents.
-This approach provides a unified framework for understanding symbol emergence, language evolution, and the nature of linguistic knowledge from the perspective of environmental adaptation and brain science.
-II-B Generative EmCom
-Numerous studies on EmCom/EmLang have been conducted.
-Major approaches are language games, MARL, and iterated learning models (ILMs) [12, 6, 49, 50, 51].
-Recently, EmCom studies based on language games, particularly referential games, have become a dominant approach in the research community of EmCom.
-These studies utilize the representation learning capability of DNNs, give agents language encoders and decoders, and make them learn to generate and interpret a complex sequence of tokens to identify its meaning.
-In a typical setting of referential games, a speaker agent encodes a given object to a sequence of tokens using a neural network, for instance, long short-term memory, and the listener agent decodes it and identifies the object among some other items.
-This approach considers EmCom as an optimization of communication protocol.
-Generally, this is in line with Shannon’s communication model [52].
-This approach focuses less on the synergy between the perceptual and communicative aspects of language.
-This approach can be regarded as a discriminative model-based approach to EmCom/EmLang.
-Unlike traditional EmCom models, generative EmCom models consider language emergence as a generative model.
-In the simplest case, a probabilistic graphical model (PGM) of generative EmCom based on CPC hypothesis can be described mathematically as follows:
-Generative model:
-$p\left(\{x^{k}\}_{k},\{z^{k}\}_{k},m\right)=p(m)\prod_{k}p\left(x^{k}\mid z^{k}\right)p\left(z^{k}\mid m\right)$ (1)
-Inference model:
-$q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)=q\left(m\mid\{z^{k}\}_{k}\right)\prod_{k}q\left(z^{k}\mid x^{k}\right)$ (2)
-where $p(\cdot)$ and $q(\cdot)$ are assumed to be original probability density functions and approximate probability functions, respectively, following the conventional notation of variational inference.
-Variables and functions used in this paper are listed in Table I.
-In this framework, $q\left(z^{k}\mid x^{k}\right)$ is interpreted as the process of representation learning performed by agent $k$, whereas $q\left(m\mid\{z^{k}\}_{k}\right)$ is hypothesized to be derived through collaborative language games among the agents.
-The overall phenomenon of symbol emergence within the group can be considered as a distributed effort to approximate $q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)$ without centralized control.
-Mathematically, the CPC frames the process of symbol emergence in terms of social representation learning, analogous to the formation of internal representations by individual brains.
-As we introduced in Section 3, the representation learning and the inference of latent variables can be naturally extended to multi-agent reinforcement learning (RL) involving EmCom.
-Additionally, the formulation of CPC from the perspective of active inference can make us understand that this generative EmCom is not only for creating communication protocol but also for the environmental adaptation as a system of agents.
-Similarly, FEP and active inference are the general explanatory theory of the brain’s adaptation to the environment.
-In summary, the framework of generative EmCom based on the CPC hypothesis offers an explanation for language evolution to serve environmental adaptation at a collective level.
-Generative EmCom offers an intriguing dual perspective, depending on whether one focuses on the collective or the individual.
-From a holistic perspective, the system appears to perform centralized representation learning of $w$ across all agents.
-Notably, $o^{k}$ can be easily extended to include multiple sensory modalities.
-From the perspective of the system overseer, or generative EmCom, multi-agent symbol emergence becomes analogous to multi-modal representation learning of external representations.
-Conversely, from the perspective of an individual agent, the system engages in multi-modal representation learning, influenced by and contributing to the inference of a shared latent variable $m$, which parameterizes the prior distribution common to all agents.
-The equivalence of these two interpretations is the cornerstone of the generative EmCom perspective.
-In essence, generative EmCom can be formulated as the representation learning of shared external representations, mediated through the representation learning of internal representations by each agent, based on multi-agent and multi-modal sensory-motor information.
-TABLE I: Nomenclature and parameter details
-$m_{t}$ Message (shared latent variable) communicated between agents at time $t$
-$o^{k}_{t}$ Optimality variable for the $k$-th agent (1: optimal, 0: not optimal)
-$z^{k}_{t}$ Internal representation (e.g., state in RL) of the $k$-th agent at time $t$
-$a^{k}_{t}$ Action of the $k$-th agent at time step $t$
-$x^{k}_{t}$ Observations (or sensory inputs) of the $k$-th agent at time $t$
-$\theta^{k}$ Global parameters of the internal models of the $k$-th agent, for instance, neural networks
-$\phi^{k}$ Parameters of language model of the $k$-th agent.
-$r^{k}$ Reward function for the $k$-th agent
-$p(\cdot)$ Original probability density function (i.e., a generative model)
-$q(\cdot)$ Approximate probability function (i.e., an inference model)
-II-C Language Game as Decentralized Bayesian Inference
-The hypothetical argument that language game can perform the decentralized Bayesian inference has a computational basis though whether actual language communication can realize such decentralized Bayesian inference in our human society is an open question.
-The MHNG is an instance of this idea.
-The MHNG comprises the following steps:
-  1. 1.
-    Perception: Speaker and listener agents (Sp and Li) observe the $d$-th object, obtain $x^{Sp}_{d}$, and $x^{Li}_{d}$ infers their internal representations $z_{d}^{Sp}$ and $z_{d}^{Li}$, respectively.
-  2. 2.
-    MH communication: Speaker mentions the name $m_{d}^{Sp}$ of the $d$-th object by sampling it from $P(m_{d}|z_{d}^{Sp},\phi^{Sp})$.
-The listener determines whether it accepts the naming with probability $\gamma={\rm min}\left(1,\frac{P(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Sp})}{P(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Li})}\right)$.
+    Signaling: Sender $S_{\boldsymbol{\phi}}$ generates a message $m$ from the observation $x$, that is, $m\sim S_{\boldsymbol{\phi}}(m|x)$.
   3. 3.
-    Learning: After MH communication was performed for every object, the listener updates its global parameters $\theta^{Li}$ and $\phi^{Li}$.
-  4. 4.
-    Turn-taking: The speaker and listener alternate their roles and go back to (1).
-It has been demonstrated that the MHNG is equivalent to the Metropolis-Hastings algorithm for inferring latent variables in a probabilistic generative model.
-This model conditions the internal representations $z^{k}$ of multiple agents, acting as representation learning machines, on a common external representation $m$.
-Although the original study assumed two agents and a categorical message $m$, the core PGM of generative EmCom underlying this theory does not make these assumptions.
-Consequently, this fundamental idea can be extended in various ways.
-As described, unlike referential games, MHNG assumes a joint attention performed by two agents.
-This assumption may seem strange from the game-theoretic approach to EmCom, such as referential games.
-However, developmental studies suggest that joint attention is fundamental to human infant language acquisition.
-Human infants develop joint attention capabilities before vocabulary explosion occurs, and joint attention serves as a crucial foundation for language acquisition [7].
-Gergely et al.
-highlighted that when children incorporate parental instructions into their learning process, they presuppose that parents have the intention to teach them [53].
-Moreover, interactions modeled by referential games—where agents identify objects and receive feedback—are rarely observed in natural language acquisition in cognitive development.
-Unlike game-theoretic approaches, MHNG, as an example of generative EmCom, assumes cooperative behavior in symbol emergence, which aligns with scholarly observations about human language acquisition [7].
-Thus, although conventional (discriminative) EmCom provides rational models for both cooperative and competitive games, the assumptions of generative EmCom better align with empirical findings in human communication behavior.
-Figure 2: (Top) Overview of the MHNG process.
-In the game, which two agents (A and B) engage in, the agents perceive a target object with joint attention and form internal representations.
-One agent (speaker) utters a sign and the other (listener) determines whether to accept it.
-Thereafter, they take turns.
-(Bottom) PGM representation, which is assumed in the MHNG.
-The MHNG is proved to be an inference process in the representation learning in a collective multi-agent system [47].
-The original idea of MH-based EmCom was introduced by Hagiwara et al.
-[48] and later generalized and formalized by Taniguchi et al.
-[47], who clarified its connection to representation learning.
-The concept of MHNG has since been extended and validated in various ways.
-Furukawa et al.
-[54] demonstrated that similar symbol emergence can be observed even when the relationship between $m$ and $z$ is head-to-head instead of tail-to-tail.
-The extension to multimodal sensory information was achieved by Hagiwara et al.
-[55], who demonstrated that MHNG can lead to symbol emergence even when agents have different sensory modalities.
-Their work also showed that modality information possessed by other agents, but not by oneself, can facilitate object category formation.
-Similar multimodal extensions have been explored with variational autoencoder (VAE)-based representation learning by Hoang et al.
-[56], though their work suggests that the mechanism for integrating multimodal sensory information influences symbol emergence.
-Although these MHNG studies involve two agents, a mathematical extension to N-agent conditions was developed by Inukai et al.
-[57], who introduced a recursive structure in communication while maintaining its characteristics as distributed Bayesian inference.
-They argued that random partner selection for MHNG can be considered as a one-sample and limited-length approximation of this approach.
-Although several generative EmCom studies define the sign $m$ as a categorical variable, this is not a necessary assumption of the framework.
-A recent study by Hoang et al.
-[58] demonstrates that sharing compositional word sequences is possible within the MHNG framework, similar to numerous EmCom studies.
-From Peirce’s semiotic perspective, signs include both compositional discrete sequences and continuous signs such as voice pitch and facial expressions.
-A recent study by You et al.
-[59] demonstrated that such continuous signs can emerge within the generative EmCom framework, whereas Saito et al.
-[60] modeled the emergence of compositional signs from continuous time-series information as signals.
-To verify the CPC hypothesis, examining whether human sign acceptance rates in joint attention naming scenarios match MHNG predictions was necessary.
-An experimental semiotic study by Okumura et al.
-[61] demonstrated that the sign acceptance rate in MHNG effectively predicts human behavior.
-MHNG serves as a basic language game to represent symbol emergence by realizing distributed Bayesian inference and inferring latent variables of the generative model, corresponding to language.
-The critical point is that the inference of the posterior distribution is performed in a decentralized manner through language games.
-This suggests that being based on the MH method is not a necessary condition.
-The MHNG-based approach represents an initial step in modeling symbol emergence (or language emergence, EmCom) as decentralized Bayesian inference.
-Future studies should explore the possibility of constructing generative EmCom models by distributing various inference methods (e.g., [62]).
-III EmCom for Multi-agent Cooperation
-III-A Generative EmCom for Multi-agent Reinforcement Learning
-Figure 3: Graphical model of generating cooperative actions for two agents.
-Communication and language are often considered to emerge to facilitate multi-agent cooperation.
-In recent years, studies on MARL with communication channels have been progressing.
-Initial methods in multi-agent deep RL include DIAL [63] and CommNet [64].
-These methods connect the networks of agents through messages, enabling the learning of necessary messages for cooperative behavior through backpropagation.
-Additionally, multi-agent deep deterministic policy gradient (MADDPG), an extension of the DDPG [65] for MARL, has been proposed [66].
-In these methods, the agents use messages sent from all other agents to learn their policies.
-Methods involving weighting or attention mechanisms have been developed to limit communication to only necessary agents [67, 68, 69, 70].
-To compute efficient messages, graph neural networks (GNNs) are used [71, 72, 73, 74, 75].
-In these studies, multiple agents are connected through a network and messages are inferred by making them differentiable variables through backpropagation.
-In other words, error information computed from the internal states of others is directly transmitted to oneself, which is an unnatural modeling from the perspective of communication among independent individuals.
-In contrast, generative EmCom allows us to formulate MARL with emergent communication in a Bayesian manner by incorporating the idea of CaI [76], which is a theory to formulate RL as a PGM [77, 78].
-Figure 3 shows a graphical model of cooperative action generation between two agents, and the details of each stochastic variable are listed in Table I.
-The behavior of each agent is generated through a Markov decision process with a prior variable $m_{t}$.
-The state $z^{k}_{t}$ of an agent at time $t$ is determined according to state $z^{k}_{t-1}$, action $a^{k}_{t-1}$, and message $m_{t}$, which is the shared latent variable:
-$z^{k}_{t}\sim p(z^{k}_{t}|m_{t},z^{k}_{t-1},a^{k}_{t-1}).$ (3)
-where $k\in\{A,B\}$ denotes an index of agents. The agent can indirectly infer the state of others through the message $m_{t}$ in a probabilistic manner.
-The optimality variable $o^{k}_{t}\in\{0,1\}$ represents the state optimality of both agents: $1$ indicates that the state and action pair are on the optimal trajectories, whereas 00 indicates it is not.
-Note that optimality is a type of probabilistic interpretation of reward functions as shown below.
-In this model, two types of optimality exist: one for each individual agent and the other for the group of agents.
-The probability $p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})$ of this optimality variable is computed using reward function $r^{k}(z^{k}_{t},a^{k}_{t})$ as follows:
-$p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})\propto\exp(r(z^{k}_{t},a^{k}_{t})).$ (4)
-For the reward of the group of agents $r^{AB}$, the following approximate reward models are learned by each agent:
-$p(o^{AB}_{t}=1|z^{k}_{t},a^{k}_{t})\approx\propto\exp(r^{AB}(z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})).$ (5)
+    Reconstruction: Receiver $R_{\boldsymbol{\theta}}$ attempts to reconstruct the original observation $x$ from the message $m$ via $R_{\boldsymbol{\theta}}(x|m)$.
+Sender $S_{\boldsymbol{\phi}}$ and receiver $R_{\boldsymbol{\theta}}$ are optimized via a gradient-based method toward successful communication.
+Conventionally, the objective function of the signaling game (to be maximized) is defined as follows [16, 106]:
+$\mathcal{J}_{\textup{MI}}(\boldsymbol{\phi},\boldsymbol{\theta})\mathrel{:=}\mathbb{E}_{p(x),S_{\boldsymbol{\phi}}(m|x)}[\log R_{\boldsymbol{\theta}}(x|m)].$ (1)
+We refer to $\mathcal{J}_{\textup{MI}}$ as the mutual information (MI)-maximizing objective function because it is known to be an evidence lower bound (ELBO) (up to constant) of the following mutual information between $X$ and $M$ [8, 99]:
+$I_{\boldsymbol{\phi}}(X;M)\mathrel{:=}\mathbb{E}_{p(x),S_{\boldsymbol{\phi}}(m|x)}\left[\log\frac{S_{\boldsymbol{\phi}}(m|x)}{\mathbb{E}_{p(x^{\prime})}[S_{\boldsymbol{\phi}}(m|x^{\prime})]}\right].$ (2)
+This implies that the conventional signaling game in the field of EmCom has been formulated as a problem of maximizing the mutual information between $X$ and $M$, where $X$ and $M$ denote random variables corresponding to the realizations $x$ and $m$, respectively.
+However, from the viewpoint of our paper, it is more fruitful to reinterpret these communication games through the lens of probabilistic generative models.
+This shift from a discriminative to a generative perspective reveals deeper connections to principles of cognition, learning, and the information-theoretic foundations of frameworks like the Variational Autoencoder (VAE) and the Information Bottleneck (IB) principle [3, 144].
+This generative viewpoint reframes the signaling game as a problem of maximizing the ELBO [130]. The objective function becomes:
 […]
-Following the theory of CaI, the optimal state sequence for both agents can be calculated by inferring state $z_{t}$ and message $m_{t}$ under the condition that the value of the optimality variables is always 1, as if the two-agent system acts as a single agent:
-$z^{A}_{t},m_{t}\sim p(z^{A}_{t},m_{t}|z^{B}_{t},o^{A}_{1:T}={\mathbf{1}},o^{AB}_{1:T}={\mathbf{1}}).$ (6)
-However, this equation has two problems: it includes the internal state $z^{B}_{t}$ of others, which cannot be observed in practice, and deriving this probability distribution analytically is difficult.
-We solve these problems by alternately inferring the following two variables:
-[…]
-$m_{1:T}\sim p(m_{1:T}|z^{A}_{1:T},z^{B}_{1:T},o^{AB}_{1:T}={\mathbf{1}}){\rm:communication.}$
-Equation (7) describes state planning, which can be computed based on the CaI framework [76].
-Equation (8) describes the inference of the message and can be formulated using the MHNG proposed by Taniguchi et al.
-[55, 47], which allows both agents to infer messages through communication without observing each other’s internal states.
-Note that MHNG enables two agents to perform sampling in (8) without simultaneous observations of $z^{A}_{t},z^{B}_{t}$.
-Thus, MHNG can be used not only for multimodal object categorization and naming but also for action coordination among multi-agents using PGM to formulate MARL, that is, modeling EmCom for multi-agent cooperation.
-III-B Generative EmCom on World models
-The generative EmCom for MARL can be extended to involve representation learning in the same way as we discussed in Section 2.3.
-This extension implies the integration of the idea of world models into generative EmCom.
-The concept of a world model represents an internal model within an agent that captures the dynamics of environmental states, their responses to the actions of the agent, and their relationships with sensory inputs [79, 18, 19].
+We refer to $\mathcal{J}_{\textup{ELBO}}$ as the ELBO-maximizing objective function, contrasting it with the MI-maximizing objective $\mathcal{J}_{\textup{MI}}$.
+By adopting the ELBO-maximizing objective function, we can introduce concepts from computational psycholinguistics into signaling games.
+To observe this, let us transform the ELBO maximizing objective function as follows:
+$\begin{split}\mathcal{J}_{\textup{ELBO}}(\boldsymbol{\phi},\boldsymbol{\theta})&\mathrel{:=}\mathbb{E}_{p(x),S_{\boldsymbol{\phi}}(m|x)}[\underbrace{\log R_{\theta}(x|m)}_{\textup{communication}}+\beta\underbrace{\log p_{\theta}(m)}_{\begin{subarray}{c}\textup{(negative)}\\ \textup{surprisal}\end{subarray}}]\\ &\phantom{\mathrel{:=}}-\beta\underbrace{\mathbb{E}_{p(x)}\mathcal{H}(S_{\phi}(M|x))}_{\begin{subarray}{c}\textup{entropy}\\ \textup{maximizer}\end{subarray}},\end{split}$ (4)
+where $\beta$ is a hyperparameter.
+Here, a term known as surprisal appears, which is a concept commonly used in computational psycholinguistics [45, 72, 66].
+Surprisal is assumed to represent the cognitive load experienced by a listener/reader (or the receiver in the signaling game) when processing a sentence.
+Therefore, the ELBO-maximizing signaling game naturally models the trade-off between information transmission and surprisal.
+A similar discussion involves modeling the trade-off between information transmission and efficiency, where studies have used the (variational) information bottleneck (IB, VIB) framework to model communication [144, 17, 127].
+In fact, VIB is proven to be a generalization of (beta-)VAE [3, 1], resulting in similar models.
+In addition, a contemporary work [128] shows that a variant framework known as a referential game can also be reformulated with an ELBO-like objective, analogous in structure to that of a conditional VAE [62, 111].
+In relation to discussions in the field of evolutionary linguistics, some studies have also incorporated ILM into the EmCom framework [103], which is another important research theme.
+The ILM is a framework that models generational changes, where supervised learning is repeatedly performed from parent agents to child agents.
+In the context of VAE and VIB, however, little discussion exists on modeling generational changes, and this remains a future challenge when considering generative symbol emergence.
+The formulation of the signaling game presented in this section, that is, generative model-based re-formulation of conventional EmCom, shares some fundamental connections with the generative frameworks discussed in later sections.
+All these formulations can be interpreted as representation learning with messages serving as latent variables in generative models when using ELBO-type (or VIB-type) objectives.
+However, these signaling game-based approaches have inherent limitations: they typically assume a simple two-agent setting with asymmetric sender-receiver roles, and their extension to populated, decentralized settings is not straightforward.
+Beyond these foundational models, the field of Semantic Communication (SC) has gained traction, aiming to transmit information based on meaning and effectiveness rather than bit-level fidelity [100].
+This has spurred research into areas such as deep learning-based semantic encoders [138], contextual reasoning for shared understanding [109], and learning goal-oriented languages from interaction [29].
+A comprehensive framework, Emergent Semantic Communication (ESC), seeks to integrate many of these threads by using causal reasoning in a neuro-symbolic architecture to create efficient emergent languages [125].
+While these approaches advance the engineering of goal-oriented communication, they do not address how a society-wide, general-purpose language emerges to reflect a collective world model, which is the central question this paper addresses.
+2.2 World Models for Individual Cognitive Agents
+The concept of a world model represents an internal model within an agent that captures the dynamics of environmental states, their responses to the actions of the agent, and their relationships with sensory inputs [37, 35, 119, 26].
 The concept of world model has its origins in the early days of AI and robotics studies [86].
 Initial studies on ML investigated techniques for agents to autonomously construct and adapt their world models [108, 116].
-Currently, the term generally refers to predictive frameworks [83], predominantly implemented using DNN architectures.
-The concept of world models is closely related to the idea of PCs.
+Currently, the term generally refers to predictive frameworks [39], predominantly implemented using deep neural network architectures.
+The concept of world models is closely related to the idea of predictive coding (PC).
 PC is the idea that the brain constantly predicts sensory information and updates its internal models to minimize prediction errors.
@@
 Learning these models enables the agent to construct a comprehensive world model that can be used for planning and decision-making in complex, partially observable environments.
-Incorporating world models into MARL based on generative EmCom is natural and reasonable.
-By estimating $z_{t}$ using the inference model of a world model $q(z_{t}|x_{1:t},a_{1:t-1})$, an agent can estimate its state.
-Agents are expected to perform EmCom in the same manner as described in Section 3.1.
-Figure 4 shows a PGM of generative EmCom on POMDPs.
-This demonstrates the theoretical connection between EmCom or language and world models.
-When considered in this way, CPC is the theory explaining that language is something that integrates world models of populated agents.
-This will be discussed in more detail in Section 5.
+This general framework has been instantiated in numerous influential models [26].
+The “World Models” approach by Ha and Schmidhuber, for instance, demonstrated that an agent could learn a compressed latent dynamics model from pixel inputs and use it to train a controller entirely within its own “dream” [38, 37].
+This line of research was significantly advanced by the Dreamer series of models, which successfully applied latent dynamics models to challenging continuous control tasks and even discrete Atari benchmarks and Minecraft [40, 39, 41, 42].
+While many of these models rely on reconstructing observations, alternative approaches have also been explored.
+For example, some models focus on learning latent dynamics through contrastive learning without reconstruction [67, 89, 90], or by incorporating specific physical priors into the latent space, such as in NewtonianVAE [54, 91].
+Furthermore, to handle information from multiple viewpoints, which is crucial in multi-agent settings, approaches like Multi-View Dreaming have been proposed to construct a world model from multi-view images using contrastive learning [63].
+A prominent recent approach to learning such world models is the Joint-Embedding Predictive Architecture (JEPA), proposed by LeCun as a key component of a pathway towards autonomous machine intelligence [70].
+Unlike generative approaches that attempt to predict missing information in pixel space, JEPA operates by predicting the representations of missing information in an abstract feature space [7].
+This non-generative, self-supervised method aims to capture the underlying dependencies of the data, encouraging the model to learn semantic and predictive features rather than superficial details.
+The JEPA framework has proven effective for learning world models from various modalities, including images (I-JEPA) [7] and videos (V-JEPA) [9].
+This line of research has recently culminated in V-JEPA 2, which extends the framework to enable robotic planning from video, further solidifying JEPA as a state-of-the-art method for an individual agent to develop its internal, predictive model of the world [6].
+In summary, all the approaches discussed in this subsection, from classic POMDP-based models to modern architectures like JEPA, are fundamentally concerned with how a single agent learns a Type 1 World Model through its own interactions.
+2.3 The Unaddressed Gap: From Individual Experience to Collective Language
+The preceding review of research in EmCom (Section 2.1) and individual world models (Section 2.2) highlights a fundamental theoretical gap.
+On one hand, studies in EmCom have explored how agents can develop shared communication protocols, but these approaches often treat language as an abstract code, detached from the rich, embodied experiences that ground meaning in the real world.
+On the other hand, computational models and general theories that could provide a comprehensive and integrative understanding of symbol and language emergence [120, 123] are still lacking.
+Specifically, these theories must address the critical interdependency between two aspects: first, the world modeling by individual agents, which is grounded in their embodiment and environmental adaptation [35, 119]; and second, the collective emergence of a language whose structure reflects this grounded world knowledge, for instance through distributional semantics [47, 36].
+In essence, a crucial question remains unaddressed: What is the theoretical and computational mechanism that bridges the subjective, internal world of the individual with the objective, external world of collective language?
+Without a clear answer, our understanding of large-scale phenomena, such as the emergence of world knowledge in LLMs, remains incomplete.
+The need for a theoretical framework that can explain the dynamic and semantic aspects of language emergence in embodied cognitive developmental systems still remains [79, 123, 13, 117], and this paper aims to provide such a framework.
+3 Generative EmCom: A Theoretical Framework
+These approaches somehow failed to construct a general framework capturing symbol emergence from the viewpoint perspective of general principles of environmental adaptation, such as the FEP, PC, and world modeling.
+Recently, world models have garnered significant attention as representation-learning models that incorporate action outputs and temporal dynamics of agent– environment interactions [38, 35, 119].
+This aligns with broader theoretical frameworks such as PC and the FEP.
+PC posits that the brain constantly predicts sensory information and updates its internal models to enhance predictability  [50], whereas FEP provides a more generalized framework explaining the self-organization of biological systems through minimization of free energy [94], which is associated with the idea of the Bayesian brain [27].
+Notably, FEP extends beyond individual cognition to explain the self-organization of cognitive and biological systems in detail [64, 33, 20], making it a promising foundation for understanding symbol emergence at both individual and collective levels.
+Figure 1: (Left) probabilistic graphical model (PGM) representation of CPC in symbol emergence, that is, EmCom or language.
+The top-down generative process corresponds to language understanding and prediction of observations, which is downward causation in the symbol emergence system.
+The bottom-up inference process corresponds to perception, representation learning, and communication, that is, language game, which is upward causation in the symbol emergence system.
+(Right) Overview of the CPC in a symbol emergence system illustrating the bidirectional process of language understanding and generation, mediated by inference through language games and representation learning.
+3.1 The CPC Hypothesis
+To address the challenges outlined in Section 2, this paper builds upon the CPC hypothesis [117, 121].
+The idea extends the principles of PC and the FEP from the individual cognitive level to a societal level [122].
+CPC posits that the emergence of symbol systems, particularly language, can be modeled as a decentralized Bayesian inference of a shared latent representation.
+It assumes that not only individual agents but also entire groups engaged in symbolic communication can be modeled as generative systems, aiming to minimize their collective free energy (CFE).
+Although PC theory suggests that individual brains constantly predict sensory information and update their internal representations, including world models.
+CPC suggests that a group of agents, for instance, a human society, predict sensory information of all of the agents and update its external representations, that is, symbol systems.
+A question is raised.
+How can we update the external representations, e.g., language, while our brains are disconnected physically?
+The CPC hypothesis suggests that a type of language game performs a decentralized Bayesian inference among the group (e.g., Taniguchi et al.
+[124], Hagiwara et al.
+[44]).
+In this framework, language games (such as naming games) can be interpreted as implementing decentralized Bayesian inference of shared representations.
+A representative example is Metropolis-Hastings Naming Game (MHNG) explained in Section 3.3.
+The CPC hypothesis argues that symbol systems emerge as a result of decentralized Bayesian inference performed collaboratively by multiple agents.
+Although the encoding of sensory information through internal representations is ensured by the plasticity of neural systems, the plasticity of external representations is guaranteed by the flexibility of our symbol systems.
+The arbitrariness of symbol systems is a widely recognized characteristic of symbols in semiotics [18].
+Peirce referred to the process by which subjects assign meaning to symbols according to culture and context as the semiosis.
+Although our brains are physically and electrically separated, they are informationally connected through communication using a flexible symbol system.
+Therefore, with appropriate communication and symbol system update algorithms, we can encode information into the symbol system as an external representation.
+In fact, the CPC hypothesis can consider that humans collectively perform this action in language emergence.
+This implies that language collectively encodes information about the world as observed by numerous agents through their sensory-motor systems.
+The CPC hypothesis study [117] did not provide a clear and detailed explanation regarding this point while proposing a new perspective on why LLMs seem to possess knowledge about the real world.
+This is one of the main topics of this paper.
+Essentially, CPC hypothesizes that human language is formed through a process of collective PC, where the symbol system emerges to maximize the predictability of multi-modal sensory-motor information obtained by members of a society, that is, minimize the CFE of a group of agents.
+This approach provides a unified framework for understanding symbol emergence, language evolution, and the nature of linguistic knowledge from the perspective of environmental adaptation and brain science.
+3.2 Formalizing Generative EmCom as Decentralized Bayesian Inference
+The CPC hypothesis posits that a group of agents, interacting with each other and their environment, can be modeled as a single, large-scale probabilistic generative model.
+This allows us to formalize the emergence of a shared symbol system, such as language, as a process of decentralized Bayesian inference.
+For a group of $K$ agents, the generative and inference processes are defined as follows:
+Generative Model: $\quad p(m,\{z_{k}\}_{k},\{x_{k}\}_{k}|\{a_{k}\}_{k})$
+$=p(m)\prod_{k=1}^{K}p(x_{k}|z_{k},a_{k})p(z_{k}|m,a_{k})$ (8)
+Inference Model: $\quad q(m,\{z_{k}\}_{k}|\{x_{k}\}_{k},\{a_{k}\}_{k})$
+$=q(m|\{z_{k}\}_{k})\prod_{k=1}^{K}q(z_{k}|x_{k},a_{k})$ (9)
+where for each agent $k$, $x_{k}$ is its observation, $a_{k}$ is its action, and $z_{k}$ is its internal representation.
+The variable $m$ represents the shared, external symbol system (i.e., message) that connects the agents.
+The inference process $q(z_{k}|x_{k},a_{k})$ corresponds to each agent’s individual representation learning, while $q(m|\{z_{k}\}_{k})$ corresponds to the collective process of symbol emergence, which can be instantiated by a language game as discussed in Section 3.3.
+Under the free energy principle, the goal of this collective system is to minimize the variational free energy (VFE), which is equivalent to maximizing the ELBO of the log-likelihood of observations.
+We term this the CFE [122, 118]:
+$F$ $=D_{\mathrm{KL}}\left[q(m,\{z_{k}\}_{k}|\{x_{k}\}_{k},\{a_{k}\}_{k})\parallel p(m,\{z_{k}\}_{k}|\{x_{k}\}_{k},\{a_{k}\}_{k})\right]$
+$=\underbrace{D_{\mathrm{KL}}\left[q(m|\{z_{k}\}_{k})\|p(m)\right]}_{\text{Collective Regularization}}$
+$\ \ \ +\sum_{k=1}^{K}\left(\underbrace{\mathbb{E}_{q}\left[-\log{p(x_{k}|z_{k},a_{k})}\right]}_{\text{Individual Prediction Error}}+\underbrace{D_{\mathrm{KL}}\left[q(z_{k}|x_{k},a_{k})\|p(z_{k}|m,a_{k})\right]}_{\text{Individual Regularization}}\right)$ (10)
+This decomposition in Eq.
+(10) is central to our framework.
+It demonstrates that minimizing a single, unified objective function, the CFE, naturally gives rise to two distinct processes.
+The second term corresponds to each agent minimizing its own individual free energy, which involves learning to accurately reconstruct its observations (the prediction error term) while keeping its internal representations regularized (the individual regularization term).
+The first term, the Collective Regularization, drives the system to form a shared symbol $m$.
+This term quantifies the cost of encoding the collective internal states $\{z_{k}\}_{k}$ into the shared message $m$.
+The emergence of language is thus framed as a process that optimizes this trade-off between individual predictive accuracy and collective communicative efficiency.
+This CFE formulation allows for a clear comparison with the ELBO-maximizing objective of signaling games presented in Section 2.1.
+While both are founded on the ELBO principle, they differ significantly in scope and structure.
+The signaling game ELBO typically models a two-agent, one-shot, and often asymmetric (sender/receiver) interaction.
+In contrast, the CFE provides a formulation for a population of $K$ agents engaged in a continuous, decentralized process of collective sense-making.
+The CFE’s explicit decomposition into individual and collective terms provides a richer, more structured model of the interplay between individual learning and social language emergence, a distinction that conventional signaling game formulations do not typically make explicit.
+Table 1: Nomenclature and parameter details
+$m_{t}$ Message (shared latent variable) communicated between agents at time $t$
+$o^{k}_{t}$ Optimality variable for the $k$-th agent (1: optimal, 0: not optimal)
+$z^{k}_{t}$ Internal representation (e.g., state in RL) of the $k$-th agent at time $t$
+$a^{k}_{t}$ Action of the $k$-th agent at time step $t$
+$x^{k}_{t}$ Observations (or sensory inputs) of the $k$-th agent at time $t$
+$\theta^{k}$ Global parameters of the internal models of the $k$-th agent, for instance, neural networks
+$\phi^{k}$ Parameters of language model of the $k$-th agent.
+$r^{k}$ Reward function for the $k$-th agent
+$p(\cdot)$ Original probability density function (i.e., a generative model)
+$q(\cdot)$ Approximate probability function (i.e., an inference model)
+3.3 The Naming Game as a Microcosm of Collective Inference
+The hypothetical argument that language game can perform the decentralized Bayesian inference has a computational basis though whether actual language communication can realize such decentralized Bayesian inference in our human society is an open question.
+The MHNG is an instance of this idea.
+The MHNG comprises the following steps:
+  1. 1.
+    Perception: Speaker and listener agents (Sp and Li) observe the $d$-th object, obtain $x^{Sp}_{d}$, and $x^{Li}_{d}$, and infer their internal representations $z_{d}^{Sp}$ and $z_{d}^{Li}$, respectively.
+  2. 2.
+    MH communication: Speaker mentions the name $m_{d}^{Sp}$ of the $d$-th object by sampling it from $p(m_{d}|z_{d}^{Sp},\phi^{Sp})$.
+The listener determines whether it accepts the naming with probability $\gamma={\rm min}\left(1,\frac{p(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Sp})}{p(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Li})}\right)$.
+  3. 3.
+    Learning: After MH communication was performed for every object, the listener updates its global parameters $\theta^{Li}$ and $\phi^{Li}$.
+  4. 4.
+    Turn-taking: The speaker and listener alternate their roles and go back to (1).
+It has been demonstrated that the MHNG is equivalent to the Metropolis-Hastings algorithm for inferring latent variables in a probabilistic generative model.
+As a result, the MHNG protocol guarantees that the CFE, defined by the KL-divergence between the agents’ collective belief and the true posterior, is monotonically non-increasing in expectation.
+This is because the underlying distributed Metropolis-Hastings sampler satisfies the detailed balance condition, a well-established result related to the Data Processing Inequality [21], ensuring convergence towards the target posterior distribution.
+This model conditions the internal representations $z^{k}$ of multiple agents, acting as representation learning machines, on a common external representation $m$.
+Although the original study assumed two agents and a categorical message $m$, the core probabilistic graphical model (PGM) of generative EmCom underlying this theory does not make these assumptions.
+Consequently, this fundamental idea can be extended in various ways.
+As described, unlike referential games, MHNG assumes a joint attention performed by two agents [^1] The assumption of joint attention is specific to the naming game, as its primary task is to establish a shared vocabulary for a common referent.
+However, the broader CPC framework does not depend on this.
+On the contrary, when applied to general multi-agent cooperation, its core strength lies in integrating different, partially observable perspectives, as we will discuss in Section 4..
+This assumption may seem strange from the game-theoretic approach to EmCom, but it is a deliberate choice made to depart from the feedback-dependent principle of success or failure that underpins many such models.
+Developmental studies suggest that such explicit feedback interactions are rarely observed in natural language acquisition and are an implausible metaphor for how infants learn [126].
+For human infants, joint attention is not a consequence of a communicative act, but rather a cognitive foundation that is in place before the vocabulary explosion occurs.
+Furthermore, children engaging in learning presuppose a pedagogical intention from adults [23], rather than simply testing a communication channel.
+Therefore, by assuming cooperative behavior grounded in joint attention, MHNG, as an example of generative EmCom, better aligns with empirical findings in human communication behavior than conventional discriminative models.
+Figure 2: (Top) Overview of the MHNG process.
+In the game, which two agents (A and B) engage in, the agents perceive a target object with joint attention and form internal representations.
+One agent (speaker) utters a sign and the other (listener) determines whether to accept it.
+Thereafter, they take turns.
+(Bottom) PGM representation, which is assumed in the MHNG.
+The MHNG is proved to be an inference process in the representation learning in a collective multi-agent system [124].
+The original idea of MH-based EmCom was introduced by Hagiwara et al.
+[44] and later generalized and formalized by Taniguchi et al.
+[124], who clarified its connection to representation learning.
+The concept of MHNG has since been extended and validated in various ways.
+The extension to multimodal sensory information was achieved by Hagiwara et al.
+[43], who demonstrated that MHNG can lead to symbol emergence even when agents have different sensory modalities.
+Their work also showed that modality information possessed by other agents, but not by oneself, can facilitate object category formation.
+Similar multimodal extensions have been explored with variational autoencoder (VAE)-based representation learning by Hoang et al.
+[48], though their work suggests that the mechanism for integrating multimodal sensory information influences symbol emergence.
+Although these MHNG studies involve two agents, a mathematical extension to $N$-agent conditions was developed by Inukai et al.
+[52], who introduced a recursive structure in communication while maintaining its characteristics as distributed Bayesian inference.
+They argued that random partner selection for MHNG can be considered as a one-sample and limited-length approximation of this approach.
+The Generative EmCom framework is general, and its core mechanisms are not limited to the exchange of simple categorical signs; they can be readily extended to handle complex, compositional languages.
+A recent study by Le Hoang et al.
+[69] demonstrates that sharing compositional word sequences is possible within the MHNG framework, similar to numerous EmCom studies.
+Furthermore, Matsui et al.
+[80] extended this concept to a full captioning game, where multiple Vision-Language Models interactively generate and refine natural language captions for images.
+From Peirce’s semiotic perspective, signs include both compositional discrete sequences and continuous signs such as voice pitch and facial expressions.
+A recent study by You et al.
+[143] demonstrated that such continuous signs can emerge within the generative EmCom framework, whereas Saito et al.
+[107] modeled the emergence of compositional signs from continuous time-series information as signals.
+To verify the CPC hypothesis, examining whether human sign acceptance rates in joint attention naming scenarios match MHNG predictions was necessary.
+An experimental semiotic study by Okumura et al.
+[92] demonstrated that the sign acceptance rate in MHNG effectively predicts human behavior.
+MHNG serves as a basic language game to represent symbol emergence by realizing distributed Bayesian inference and inferring latent variables of the generative model, corresponding to language.
+The critical point is that the inference of the posterior distribution is performed in a decentralized manner through language games.
+This suggests that being based on the MH method is not a necessary condition.
+The MHNG-based approach represents an initial step in modeling symbol emergence (or language emergence, EmCom) as decentralized Bayesian inference.
+Future studies should explore the possibility of constructing generative EmCom models by distributing various inference methods (e.g., Hoang et al.
+[49]).
+4 Emergence of a Collective World Model through Multi-Agent Interaction and Cooperation
+The previous section used the MHNG to illustrate decentralized inference under the simplified assumption of joint attention.
+This section expands upon that principle, turning to more general multi-agent systems where joint attention is not required.
+Here, the essence of the CPC framework becomes clear: to integrate the diverse, partially observable perspectives of multiple agents through communication.
+We will explore how an emergent language allows agents to resolve their individual uncertainties and coordinate their actions by inferring a shared understanding of the world.
+This exploration begins with the application of Generative EmCom to MARL, considering that Type 1 world models are theoretically rooted in reinforcement learning studies.
+4.1 From Communication to Cooperation in Multi-Agent Systems
+Figure 3: Graphical model of generating cooperative actions for two agents.
+Communication and language are often considered to emerge to facilitate multi-agent cooperation.
+In recent years, studies on MARL with communication channels have been progressing.
+Initial methods in multi-agent deep RL include DIAL [30] and CommNet [115].
+These methods connect the networks of agents through messages, enabling the learning of necessary messages for cooperative behavior through backpropagation.
+Additionally, multi-agent deep deterministic policy gradient (MADDPG), an extension of the DDPG [75] for MARL, has been proposed [78].
+In these methods, the agents use messages sent from all other agents to learn their policies.
+Methods involving weighting or attention mechanisms have been developed to limit communication to only necessary agents [59, 56, 60, 53].
+To compute efficient messages, graph neural networks (GNNs) are used [19, 76, 2, 87, 101].
+In these studies, multiple agents are connected through a network and messages are inferred by making them differentiable variables through backpropagation.
+In other words, error information computed from the internal states of others is directly transmitted to oneself, which is an unnatural modeling from the perspective of communication among independent individuals.
+In contrast, generative EmCom allows us to formulate MARL with EmCom in a Bayesian manner by incorporating the idea of CaI [71], which is a theory to formulate RL as a PGM [28, 85].
+Figure 3 shows a graphical model of cooperative action generation between two agents, and the details of each stochastic variable are listed in Table 1.
+The behavior of each agent is generated through a Markov decision process with a prior variable $m_{t}$.
+The state $z^{k}_{t}$ of an agent at time $t$ is determined according to state $z^{k}_{t-1}$, action $a^{k}_{t-1}$, and message $m_{t}$, which is the shared latent variable:
+$z^{k}_{t}\sim p(z^{k}_{t}|m_{t},z^{k}_{t-1},a^{k}_{t-1}).$ (11)
+where $k\in\{A,B\}$ denotes an index of agents. The agent can indirectly infer the state of others through the message $m_{t}$ in a probabilistic manner.
+The optimality variable $o^{k}_{t}\in\{0,1\}$ represents the state optimality of both agents: $1$ indicates that the state and action pair are on the optimal trajectories, whereas 00 indicates it is not.
+Note that optimality is a type of probabilistic interpretation of reward functions as shown below.
+In this model, two types of optimality exist: one for each individual agent and the other for the group of agents.
+The probability $p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})$ of this optimality variable is computed using reward function $r^{k}(z^{k}_{t},a^{k}_{t})$ as follows:
+$p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})\propto\exp(r(z^{k}_{t},a^{k}_{t})).$ (12)
+For the group of agents, the true optimality $o^{AB}_{t}$ depends on the joint state and actions, with its probability being proportional to a global reward function $r^{AB}$:
+$p(o^{AB}_{t}=1|z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})\propto\exp(r^{AB}(z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})).$ (13)
+However, since each agent $k$ cannot observe others’ internal variables, it must learn an approximate model of this group optimality conditioned only on its own state $z^{k}_{t}$ and action $a^{k}_{t}$.
+Agent $k$ therefore learns its own predictive model of group success as:
+$p(o^{AB}_{t}=1|z^{k}_{t},a^{k}_{t})\propto\exp(\hat{r}^{AB}_{k}(z^{k}_{t},a^{k}_{t})),$ (14)
+where $\hat{r}^{AB}_{k}(z^{k}_{t},a^{k}_{t})$ is a function learned by agent $k$ that locally approximates the global reward.
+This formulation eliminates the necessity for an agent to access others’ internal variables to model group success, while still allowing it to contribute to the collective goal.
+Following the theory of CaI, the optimal state sequence for both agents can be calculated by inferring state $z_{t}$ and message $m_{t}$ under the condition that the value of the optimality variables is always 1, as if the two-agent system acts as a single agent:
+$z^{A}_{t},m_{t}\sim p(z^{A}_{t},m_{t}|z^{B}_{t},o^{A}_{1:T}={\mathbf{1}},o^{AB}_{1:T}={\mathbf{1}}).$ (15)
+However, this equation has two problems: it includes the internal state $z^{B}_{t}$ of others, which cannot be observed in practice, and deriving this probability distribution analytically is difficult.
+We solve these problems by alternately inferring the following two variables:
+[…]
+$m_{1:T}\sim p(m_{1:T}|z^{A}_{1:T},z^{B}_{1:T},o^{AB}_{1:T}={\mathbf{1}}){\rm:communication.}$
+Equation (16) describes state planning, which can be computed based on the CaI framework [71].
+Equation (17) describes the inference of the message and can be formulated using the MHNG proposed by Hagiwara et al.
+[43] and Taniguchi et al.
+[124], which allows both agents to infer messages through communication without observing each other’s internal states.
+Note that MHNG enables two agents to perform sampling in (17) without simultaneous observations of $z^{A}_{t},z^{B}_{t}$.
+Thus, MHNG can be used not only for multimodal object categorization and naming but also for action coordination among multi-agents using PGM to formulate MARL, that is, modeling EmCom for multi-agent cooperation.
+This theoretical connection has been recently instantiated in the MARL-CPC framework, which successfully applies the principles of CPC to on-policy RL algorithms and demonstrates effective communication in non-cooperative settings [141].
 Figure 4: Graphical model of generative EmCom involving world models.
-Several studies linking EmCom and world models have been conducted.
-Recent studies on EmLang have focused on MARL approaches, aiming to autonomously develop languages to solve tasks [2].
-Because each agent typically has access only to its own observations, inferring the underlying state of the environment becomes crucial.
-Consequently, researchers have explored enabling RL agents to develop EmLangs for inter-agent communication while simultaneously learning world models to infer environmental states.
-Notably, not all MARL communication studies focus on developing symbol-based languages; some treat communication as simple information exchange.
-This distinction separates “ EmLang” from “learning tasks with communication.” [49]
-Alexander et al.
-proposed a message-conditioned world model that improves the quality of communication among agents, facilitating the emergence of meaningful messages and enhancing the interpretability of the model [84].
-Similarly, Kenzo Lobos-Tsunekawa et al.
-extended the widely used Dreamer world model by introducing a global world model, used exclusively during training.
-By incorporating information from all agents, they achieved improved sample efficiency through shared imagination [85].
-Toledo et al.
-also used Dreamer but introduced GNNs to facilitate communication; however, their work is classified under ”learning tasks with communication” [86] because it does not aim to develop a model of EmLang.
-IV Relationship with Conventional EmCom
-Recent years have witnessed an increasing presence of research papers focused on EmCom at ML conferences such as NeurIPS and ICLR.
-This trend underscores the renewed interest in EmCom.
-In the field of EmCom, Lewis’ signaling game [87] framework is often used as a model of communication while other various formulations have also been proposed [35, 36, 37].
-The game players, or agents, are optimized to achieve some goals, resulting in the communication protocols, which are referred to as EmLang.
-In the EmCom field, the extent to which EmLangs resemble human languages is frequently questioned from the perspectives of evolutionary and/or computational linguistics.
-For instance, questions have been raised about whether EmLangs have compositionality [88, 89] or whether they follow the well-known statistical universal properties of natural languages [90, 91, 92, 93].
-Additionally, cases exist where cognitive constraints were modeled to induce characteristics of natural languages [94, 95].
-The signaling game is a simple communication model that involves only a sender $S_{\boldsymbol{\phi}}(m|x)$ and a receiver $R_{\boldsymbol{\theta}}(x|m)$ and only allows unidirectional communication from the sender to the receiver.
-At each play, the signaling game proceeds as follows:
-  1. 1.
-    Observation: Sender $S_{\boldsymbol{\phi}}$ obtains an observation $x$, that is, $x\sim P(x)$.
-  2. 2.
-    Signaling: Sender $S_{\boldsymbol{\phi}}$ generates a message $m$ from the observation $x$, that is, $m\sim S_{\boldsymbol{\phi}}(m|x)$.
-  3. 3.
-    Reconstruction: Receiver $R_{\boldsymbol{\theta}}$ attempts to reconstruct the original observation $x$ from the message $m$ via $R_{\boldsymbol{\theta}}(x|m)$.
-Sender $S_{\boldsymbol{\phi}}$ and receiver $R_{\boldsymbol{\theta}}$ are optimized via a gradient-based method toward successful communication.
-Conventionally, the objective function of the signaling game (to be maximized) is defined as follows [96, 97]:
-$\mathcal{J}_{\textup{MI}}(\boldsymbol{\phi},\boldsymbol{\theta})\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}[\log R_{\boldsymbol{\theta}}(x|m)].$ (12)
-We refer to $\mathcal{J}_{\textup{MI}}$ as the MI-maximizing objective function because it is known to be a variational lower bound (up to constant) of the following mutual information between $X$ and $M$ [98, 99]:
-$I_{\boldsymbol{\phi}}(X;M)\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}\left[\log\frac{S_{\boldsymbol{\phi}}(m|x)}{\mathbb{E}_{P(x^{\prime})}[S_{\boldsymbol{\phi}}(m|x^{\prime})]}\right].$ (13)
-This implies that the conventional signaling game in the field of EmCom has been formulated as a problem of maximizing the mutual information between $X$ and $M$, where $X$ and $M$ denote random variables corresponding to the realizations $x$ and $m$, respectively.
-In contrast, a recent study [100] has proposed a reformulation of the signaling game as a problem of maximizing the evidence lower bound (ELBO):
-[…]
-We refer to $\mathcal{J}_{\textup{ELBO}}$ as the ELBO-maximizing objective function, contrasting it with the MI-maximizing objective $\mathcal{J}_{\textup{MI}}$.
-By adopting the ELBO-maximizing objective function, we can introduce concepts from computational psycholinguistics into signaling games.
-To observe this, let us transform the ELBO maximizing objective function as follows:
-$\begin{split}\mathcal{J}_{\textup{ELBO}}(\boldsymbol{\phi},\boldsymbol{\theta})&\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}[\underbrace{\log R_{\theta}(x|m)}_{\textup{communication}}+\beta\underbrace{\log P_{\theta}(m)}_{\begin{subarray}{c}\textup{(negative)}\\ \textup{surprisal}\end{subarray}}]\\ &\phantom{\mathrel{:=}}-\beta\underbrace{\mathbb{E}_{P(x)}\mathcal{H}(S_{\phi}(M|x))}_{\begin{subarray}{c}\textup{entropy}\\ \textup{maximizer}\end{subarray}},\end{split}$ (15)
-Here, a term known as surprisal appears, which is a concept commonly used in computational psycholinguistics [101, 102, 103].
-Surprisal is assumed to represent the cognitive load experienced by a listener/reader (or the receiver in the signaling game) when processing a sentence.
-Therefore, the ELBO-maximizing signaling game naturally models the trade-off between information transmission and surprisal.
-A similar discussion involves modeling the trade-off between information transmission and efficiency, where studies have used the (variational) information bottleneck (IB, VIB) framework to model communication [104, 105, 106].
-In fact, VIB is proven to be a generalization of (beta-)VAE [107, 108], resulting in similar models.
-In addition, a contemporary work [109] shows that a variant framework known as a referential game can also be reformulated with an ELBO-like objective, analogous in structure to that of a conditional VAE [110, 111].
-In relation to discussions in the field of evolutionary linguistics, some studies have also incorporated ILM into the EmCom framework [112], which is another important research theme.
-The ILM is a framework that models generational changes, where supervised learning is repeatedly performed from parent agents to child agents.
-In the context of VAE and VIB, however, little discussion exists on modeling generational changes, and this remains a future challenge when considering generative symbol emergence.
-The formulation of the signaling game presented in this section, that is, generative model-based re-formulation of conventional EmCom, shares some fundamental connections with the generative EmCom and CPC frameworks discussed in other sections.
-All these formulations can be interpreted as representation learning with messages (signals, signs, or sentences) serving as latent variables in generative models when using ELBO-type (or VIB-type) objectives.
-However, notable differences exist: Although the CPC hypothesis does not specify a particular method for posterior distribution estimation, the signaling game explicitly adopts amortized variational inference.
-Furthermore, the signaling game assumes asymmetric roles between senders and receivers, which is not necessarily the case in the CPC frameworks.
-Additionally, the extension of ELBO-type formulation to populated settings, where multiple agents interact, remains an open challenge in the signaling game framework and requires further investigation [^1] An ELBO-type formulation based on the FEP for populated settings of generative EmCom is described in [45].
-However, an inference procedure, that is, a language game, for enabling agents to let a symbol system emerge has not been described..
-V Large Language Model as Collective World models
+4.2 Language as an External Representation for Integrating World Models
+As established in Section 2.2, an individual agent’s world model is an internal mechanism used to predict future latent states based on current states and actions, i.e., it learns the state transition dynamics $p(z_{t+1}|z_{t},a_{t})$.
+In a multi-agent setting where each agent has only partial observations of the environment, each agent learns its own subjective world model.
+Without a mechanism for exchanging information, these individual world models remain isolated, and the agents lack a way to form a shared, consistent understanding of the global environment.
+Our Generative EmCom framework introduces such a mechanism by incorporating a shared latent variable, the message $m$, into the world model of each agent.
+This fundamentally alters the state transition dynamics.
+As depicted in the PGM for a single agent with communication (Figure 4), the world model of agent $k$, parameterized by $\theta^{k}$, now learns the dynamics conditioned on the message: $p(z^{k}_{t+1}|z^{k}_{t},a^{k}_{t},m_{t})$.
+Here, the message $m_{t}$ is not part of the action space; instead, it acts as a conditioning variable that directly influences the agent’s prediction of its future internal state $z^{k}_{t+1}$.
+This latent state, in turn, informs the agent’s policy, $\pi^{k}$, which generates the subsequent action.
+This architecture creates a powerful feedback loop for knowledge integration.
+As formalized in our framework (Section 3.2), the message $m_{t}$ is inferred from the collective internal states of all agents, $\{z^{k}_{t}\}_{k}$.
+In turn, this collectively-informed message influences the future internal state of each individual agent.
+Therefore, the process of learning to use the language $m$ effectively becomes a process of learning to align these individual world models.
+Each agent learns to encode its own unique knowledge into the shared language and to decode the language to update its own world model with information it could not observe directly.
+This perspective aligns with recent works that have demonstrated the effectiveness of message-conditioned world models for improving coordination [22, 77, 88].
+Through this process, the emergent language $m$ becomes more than just a set of signals for coordinating immediate actions.
+It becomes the shared, external representation that embodies the integrated knowledge of the individual world models.
+It is the medium through which a collective of agents, each with a subjective Type 1 World Model, can construct a shared, objective Type 2 World Model (i.e., Model of the World), allowing the group to model the world from a unified perspective that transcends any single agent’s limited view.
 Figure 5: Three levels of the relationship between language, perceptual, and action information.
@@
 Right: probabilistic generative models corresponding to each of (A) – (C).
-V-A Our argument
-Based on the discussion in Sections 2 and 3, we explain the main argument of this paper, stating that LLMs are collective world models.
-To clarify the link between this argument and existing studies relating language and perceptual information, Figure 5 shows three levels of tasks relating language to visual information (an example of multimodal sensory information) and action streams, along with PGMs for each.
-These represent (A) image captioning and generation tasks, corresponding language to a still image [113, 114, 115]; (B) video captioning and generation tasks, corresponding language to a video, that is, a sequence of visual stimuli [116, 117]; and (C) action-dependent video captioning and generation, corresponding language to dynamic perceptual and action information, respectively [^2] Note that we redefine the generative process of world models assuming that the policies of agents that generate $a_{t}$ from $z_{t}$ are also a part of the generative process of this world model..
-In particular, from the perspective of representation learning, (A) and (B) correspond to representation learning of images and videos, whereas (C) corresponds to world models explained in Section III-B.
-Language-conditioned world models and robotics foundation models, known as vision-language-action models, studied in robotics and autonomous vehicles correspond to (C) [118, 119, 120, 121, 122].
-Corresponding to the three levels of captioning and generation tasks, we can consider generative EmCom.
-Figure 5 describes PGMs for generative EmCom corresponding to the three levels, respectively.
-Notably, they are only an instance of the generative model shown in the CPC hypothesis (Figure 1) and generalization of studies based on the MHNG described in Sections 2 and 3.
 Figure 6: PGMs for generative EmCom corresponding to three levels of complexity: (A) image-, (B) video-, and (C) action- and video-based tasks.
 These models represent instances of the CPC hypothesis and generalizations of MHNG studies.
-Here, $\theta^{k}$ represents the global parameters of the $k$-th agent, for instance, the parameters of neural networks for representation learning.
-In (C), the global parameter $\bar{\theta}^{k}$ includes not only the parameters of the prediction model of the world (i.e., Equations 9 and 10 ) but also those of policies $p(a_{t}|z_{t})$.
-Let us consider (C) as a generative model corresponding to a single super-agent, virtually.
-The super-agent has multimodal sensory information combining the observations $\{x^{k}_{1:t}\}_{k}$ and actions $\{a^{k}_{1:t}\}_{k}$ of every agent.
-In this case, (C) can be considered as a type of hierarchical world model combining $K$ agents and integrating their experiences.
-We can refer to the world model owned by the super-agent as a collective world model, where $\{\phi,\{\theta^{k}\}_{k}\}$ are its parameters.
-From this perspective, the corpus data that comprise sentences uttered by agents are considered as a sample from approximate posterior distributions $m_{j}^{[i]}\sim q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})$, where $m_{j}^{[i]}$ denotes the $i$-th sentence describing the $j$-th observation $\{x^{k}_{j,1:t}\}_{k}$ and action $\{a^{k}_{j,1:t}\}_{k}$.
-Language modeling implies approximating the distribution of word sequences $p(m|\phi^{k})$ in a corpus, where $\phi^{k}$ denotes a parameter of the language model of the $k$-th agent.
-This implies that the language model approximates the posterior distribution $q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})\approx p(m|\theta^{LM})$.
-The rationale behind distributional semantics as modeled by LLMs can be comprehended through the CPC framework.
-When language emerges through CPC, it functions as a latent space that develops through representation learning based on the collective experiences of all agents — in this case, humans.
-These experiences encompass not only external sensory information but also interception and behavioral decision-making.
-Within this framework, sentences can be mathematically considered as samples drawn from the posterior distribution over messages $m$, formally expressed as $q(m|\{x^{k},a^{k}\}_{k})$, where $x^{k}$ and $a^{k}$ represent the observations and actions of agent $k$, respectively.
-This formulation naturally leads to the structural reflection of the distribution of $\{x^{k},a^{k}\}_{k}$ in that of $m$, analogous to the working mechanism of representation learning in individual agents.
-A key insight is that language serves as a shared external representation that effectively combines the world models of multiple human agents.
-When examining the graphical model depicted in Figure 6 (C) as a cognitive model of a single ”super-agent,” we can interpret it as a factorized hierarchical world model incorporating $K$ distinct modalities.
-Through this lens, an LLM that models the distribution $q(m|\{x^{k},a^{k}\}_{k})$ can be considered as a collective world model — one that integrates the diverse experiences and knowledge of multiple agents into a unified representational space.
-This provides a theoretical foundation for understanding why LLMs trained on large-scale human-generated text corpora appear to capture rich world knowledge.
-The emergence of these phenomena can be comprehensively understood through the lens of the CPC hypothesis.
-The hypothesis suggests that language evolution occurs through a process of CPC, where the symbol system (language) emerges to maximize the predictability of multimodal sensory-motor information obtained by members of a society.
-This theoretical framework helps explain why LLMs, trained on the linguistic outputs of this process, inherit structural properties that reflect collective human knowledge and experience.
-The CPC hypothesis thus provides a principled explanation for the acquisition of world knowledge by language models without direct sensorimotor experience by tapping into the accumulated wisdom encoded in human language through this collective process.
-V-B Existing argument
-LLMs learn to predict text tokens in an autoregressive manner, and unlike human cognition or prediction, they are not designed to interact with environments or achieve specific goals beyond word prediction.
-Nevertheless, some studies have suggested that LLMs may implicitly contain a world model that simulates our physical world.
-Hao et al.
-proposed “reasoning via planning,” where an LLM is reused as a world model to perform logical reasoning, employing Monte Carlo tree search to reason in a more planned manner [123].
-This method demonstrated higher performance compared to the existing chain-of-thought approach.
-Li et al.
-used a GPT model trained only on Othello game logs and confirmed that the model made predictions based on an internal world model by performing tests using explicit intervention operations that changed the activation of the middle layer of the model [124].
-This suggests that LLM does not simply rely on statistical inference but constructs some degree of world model in systems such as games.
-In contrast, Wang et al.
-developed a benchmark involving text games to test whether transition models or policies can be derived from language models and revealed that LLMs are not reliable world simulators [125].
-Although these studies investigate whether LLMs can be used as world models, studies examining whether LLMs inherently possess such models are underway.
-Gurnee et al.
-demonstrated that LLMs acquire spatiotemporal linear representations of the real world, specifically identifying individual spatial and temporal neurons [21].
-The authors argue that this corresponds to the basic elements of a world model, although they did not show the acquisition of a causal model that includes actions.
-Andreas argues that during the process of predicting the next token, a language model infers the internal states, such as the beliefs, desires, and intentions of the speaker, from the context and reflects them in text generation or speech directed at other agents, suggesting the presence of a world model within the linguistic space [126].
-Previous studies also indicated the possibility of representing belief states and intentions during text modeling.
-However, this pertains to agents within the linguistic space and does not explain connections with data derived from the physical world.
-[…]
-VI Conclusion
+4.3 Collective World Model as an Abstractive Prior over Embodied Experiences
+To understand how language integrates individual world models into a collective one, we first consider the hierarchy of experiences that language can represent, as illustrated in Figure 5.
+This hierarchy progresses through three levels of increasing complexity.
+Level A represents the bidirectional relationship between a static observation, such as an image $x_{k}$, and a linguistic description $m$.
+This includes both image captioning (inferring $m$ from $x_{k}$) and image generation (generating $x_{k}$ from $m$[134, 139, 102].
+Level B extends this to dynamic sequences, where the language $m$ corresponds to a stream of observations over time, $\{x^{k}_{t}\}_{t=1:T}$, as seen in video captioning and generation tasks [132, 140].
+The hierarchy culminates in Level C, which models the full sensorimotor loop of an embodied agent.
+Here, language $m$ is grounded in both dynamic perception $\{x^{k}_{t}\}_{t=1:T}$ and a sequence of actions $\{a^{k}_{t}\}_{t=1:T}$.
+This final level corresponds to the world models used in modern robotics, often called Vision-Language-Action (VLA) models, which learn from an agent’s own embodied interactions [5, 61, 25, 145, 58].
+Crucially, in a collective setting, the shared message $m$ is inferred from the observations of all $K$ agents, e.g., $m\sim q(m|\{x^{k}_{1:T}\}_{k})$.
+This collective inference process itself constitutes a form of (external) representation learning, where the structure of the shared language $m$ comes to reflect the integrated knowledge of the entire group.
+The generative process for each of these three levels can be explicitly formulated as a Probabilistic Graphical Model (PGM), as shown in Figure 6.
+At Level A, the PGM captures the relationship between a static image observation $x^{k}$ and a message $m$, mediated by an internal representation $z^{k}$.
+At Level B, this model is extended to handle temporal dynamics, where a sequence of video frames $\{x^{k}_{1:T}\}$ informs a sequence of internal states $\{z^{k}_{1:T}\}$.
+As shown in the PGMs for Level A and Level B (Figure 6 (A) and (B)), the processes are governed by two sets of parameters for each agent $k$: $\theta^{k}$ represents the parameters of the agent’s internal model (i.e., the VLM), while $\phi^{k}$ denotes the parameters of the prior over the shared message $m$.
+Finally, Level C depicts the complete generative model for an embodied agent, where both sensory observations $\{x^{k}_{1:T}\}$ and actions $\{a^{k}_{1:T}\}$ jointly determine the agent’s internal state trajectory.
+The parameters of this comprehensive model, $\bar{\theta}^{k}$, therefore encompass not only the agent’s predictive world model but also its policy.
+These PGMs represent concrete instances of the overarching CPC hypothesis.
+This principle generalizes directly as we move up the hierarchy to a collective of fully embodied VLA agents (Level C).
+Here, the framework models a population of agents, each possessing its own sophisticated world model for navigating its environment (Figure 6 (C)).
+The interactive, decentralized process of these agents developing a shared language to communicate about their sensorimotor experiences ($(\{x^{k}_{1:T}\}_{k},\{a^{k}_{1:T}\}_{k})$ becomes mathematically equivalent to the process of learning a single, abstract, hierarchical world model that spans the entire group.
+Therefore, we argue that embodied symbol emergence, under the CPC framework, is collective world modeling.
+The emergent language $m$ is not merely a tool for this process; it is the very instantiation of the Collective World Model.
+It functions as a highly abstract structured prior over the entire space of possible embodied experiences within the society.
+The dynamic, interactive process of collective sense-making described above does not simply vanish; it leaves behind a static artifact in the form of a text corpus.
+We can think of this corpus as a “fossil record” of a society’s ongoing, collective inference about the world [^2] This idea has also been applied to model the process of scientific activity itself from the perspective of CPC, a framework known as CPC as a model of science (CPC-MS) [122]..
+From this perspective, each sentence or document within a corpus can be viewed as a sample from an underlying posterior distribution.
+More formally, we consider a sentence $m_{j}^{[i]}$ describing a particular event or situation $j$ to be a sample from the approximate posterior $q(m|\{\{x^{k}_{t}\}_{t=1:T},\{a^{k}_{t}\}_{t=1:T}\}_{k})$.
+It is this product, rich with the implicitly encoded structure of the collective world model, that serves as the training data for LLMs.
+5 LLM as Collective World models
+5.1 Language Corpora as Samples from the Collective World Model
+The preceding sections have established our central premise: that human language, as a product of CPC, can be understood as a Collective World Model externalized in a shared symbolic system.
+Building on this foundation, we now connect this theory to the training of LLMs.
+As argued in Section 4.3, each sentence $m$ within a text corpus is not an isolated artifact but can be viewed as a sample drawn from a complex posterior distribution, conditioned on the unobserved, collective sensorimotor experiences of the society that generated it:
+$m\sim q(m|\{\{x^{k}_{t}\}_{t=1:T},\{a^{k}_{t}\}_{t=1:T}\}_{k})$ (18)
+The vast web-scale corpora used to train modern LLMs, therefore, represent a massive dataset of samples drawn from this collective, world-grounded inference process.
+The fundamental objective of an LLM during pre-training, typically next-token prediction, is an autoregressive method to learn a model, which we denote as $p_{\text{LM}}(m)$, that approximates the marginal distribution $p(m)$ of these textual samples.
+From this perspective, the task of an LLM is not merely to learn statistical correlations in text, but to implicitly model the output of the complex, embodied, and collective process that produced the text in the first place.
+5.2 The LLM as a Reconstructor of Collective Representation
+As established in the previous section, LLMs are trained to model the distribution $P(m)$ of language corpora, which are themselves samples from a collective, world-grounded process.
+When a trained LLM processes a given sentence $m$, we can posit that it forms a corresponding high-level internal representation or latent state, which we denote as $z_{\text{LM}}$.
+This constitutes an inference or recognition process within the LLM: $q_{\text{LM}}(z_{\text{LM}}|m)$.
+This allows us to frame the end-to-end process—from human experience to LLM representation—as a two-stage process of representation transformation, as illustrated in Figure 7.
+  1. 1.
+    Stage 1: Generation by Human Society (Encoding): The collective internal states of all agents $\{z^{k}\}_{k}$, which are grounded in sensorimotor experience, are compressed and encoded into the discrete, symbolic form of language, $m$.
+  2. 2.
+    Stage 2: Reconstruction by LLM (Inference): The LLM, having learned the statistical structure of language, takes a sentence $m$ as input and infers its own internal, continuous representation $z_{\text{LM}}$.
+This can be seen as a reconstruction of a high-dimensional representation from the symbolic code.
+Figure 7: A schematic of the representation reconstruction process.
+Human society acts as a large-scale encoder, transforming embodied experiences (Type 1 World Models) from sensorimotor interactions into an externalized representation, i.e., language ($m$).
+The LLM then acts as a decoder or reconstructor, inferring its own internal latent state ($z_{\text{LM}}$) from language.
+This process allows the LLM to learn an internal model (Type 2 World Model) whose structure mirrors the original collective world model.
+The crucial implication of this two-stage process, $\{z^{k}\}_{k}\rightarrow m\rightarrow z_{\text{LM}}$, is the resulting structural alignment [121].
+The language $m$ is the sole informational bottleneck between the collective human mind and the LLM.
+Therefore, for an LLM to effectively model the distribution of $m$, it must develop an internal latent space $z_{\text{LM}}$ whose relational structure necessarily mirrors the relational structure of the original source space $\{z^{k}\}_{k}$.
+While the LLM does not reconstruct the specific values of $\{z^{k}\}_{k}$, it learns to reconstruct the geometry of the conceptual space that generated the language.
+This provides a powerful mechanism for transferring world structure from a society of embodied agents to a disembodied language model.
+5.3 Explaining Distributional Semantics and Representational Alignment
+This framework of structural representation reconstruction provides a principled explanation for the otherwise mysterious emergence of world knowledge in LLMs.
+For decades, it has been observed that language models capture the relational structure of the world, a phenomenon known as distributional semantics [47, 79].
+Even before modern LLMs, models like word2vec could perform analogical reasoning, such as computing "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [81, 82], implying that the geometry of their embedding spaces reflects real-world conceptual relationships.
+Our framework explains this as a natural consequence of the representation reconstruction process.
+To be precise, the reason the statistical relationships between words in an LLM’s latent space ($z_{\text{LM}}$) reflect the structure of the real world is because they are reconstructing the structure of the collective human internal representations ($\{z^{k}\}_{k}$).
+It is this set of internal representations that, as a whole, constitutes the collective model used by the society to predict the world.
+The LLM, in learning to model the distribution of $m$, naturally develops an internal latent space that mirrors the relational structure of the collective human representations that originally encoded that world knowledge.
+The structure is not learned from a vacuum; it is inherited.
+This framework of representation reconstruction provides a principled explanation for the otherwise mysterious emergence of world knowledge in LLMs.
+For decades, it has been observed that language models capture the relational structure of the world, a phenomenon known as distributional semantics [47, 79].
+Even before modern LLMs, models like word2vec could perform analogical reasoning, such as computing "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [81, 82], implying that the geometry of their embedding spaces reflects real-world conceptual relationships.
+Our framework explains this as a natural consequence of the representation reconstruction process.
+The reason the statistical relationships between words in an LLM’s latent space ($z_{\text{LM}}$) reflect the structure of the real world is that the language ($m$) it was trained on was generated by a society of agents collectively trying to predict that world.
+The LLM, in learning to model the distribution of $m$, inevitably develops an internal latent space that reconstructs the relational structure of the collective human representations ($\{z^{k}\}_{k}$) that originally encoded that world knowledge.
+The structure is not learned from a vacuum; it is inherited.
+This process is how an LLM acquires a powerful Model of the World (Type 2) by statistically analyzing the linguistic output of a society of agents, each possessing their own subjective World Model as an Internal Model (Type 1).
+This perspective also helps situate our hypothesis within the broader landscape of research on world models in LLMs.
+Several studies have compellingly demonstrated that LLMs can be used as world models for specific tasks, for example in planning [46] or for modeling game states [74].
+While these works establish that world-like representations can emerge, they primarily focus on how these representations can be leveraged.
+Our hypothesis, in contrast, provides a more fundamental explanation for why such rich, world-congruent knowledge is present in the first place.
+Similarly, Andreas argues that LLMs model the beliefs and intentions of agents within the linguistic space [4].
+Our framework complements and grounds this view by proposing that these linguistic agent models emerge precisely because the language itself is a reflection of a collective model formed by embodied agents interacting within a shared physical environment.
+Finally, our hypothesis offers a particularly parsimonious explanation for the representational alignment observed between different modalities.
+Huh et al.
+introduced the platonic representation hypothesis, proposing that internal representations learned from language ($q_{L}(z|m)$) and vision ($q_{V}(z|x)$) converge toward a similar latent structure [51].
+Within the CPC framework, this alignment is not a surprising outcome that requires a pre-existing platonic ideal, but an expected one.
+Since human language ($m$) emerges as a structured representation of collective observations of the world ($x=\{x^{k}\}_{k}$), it is natural that a model trained on language ($q_{L}(z|m)$) would learn a latent space whose structure mirrors that of a model trained on direct observation ($q_{V}(z|x)$).
+Our framework thus provides a generative and developmental explanation for this empirically observed alignment, rooted in the principles of collective inference.
+6 Discussion
 This study proposed a theoretical framework that unified EmCom, world models, and LLMs through the lens of CPC.
@@
 We showed the application of this framework to MARL and provided a novel perspective on LLMs as collective world models that integrate the diverse experiences and knowledge of multiple agents into a unified representational space.
+6.1 Implications for Artificial Intelligence and Cognitive Science
 The theoretical framework proposed here offers several important insights.
@@
 Third, it suggests that the success of LLMs may be fundamentally linked to their ability to capture and integrate the collective world models acquired by humans, i.e., through embodied sensorimotor interactions with the world.
+6.2 Assumptions and Limitations of the Proposed Framework
 However, several limitations of the current work should be addressed.
@@
 Third, the current formulation may not fully capture the dynamic, interactive nature of human language evolution, for instance, language evolution over generations and through the interaction of several agents in an open world.
+6.3 Future Directions
 From the perspective of world models, understanding the influence of collective world models on the environmental adaptation of agents is also important.
@@
 Our argument and theoretical framework initiate several promising directions for future studies.
-Although we have established the basic connections between emergent communication, world models, and LLMs, significant work remains to validate and extend these ideas.
+Although we have established the basic connections between EmCom, world models, and LLMs, significant work remains to validate and extend these ideas.
 Key priorities include developing experimental paradigms to test the CPC hypothesis, creating more sophisticated computational models and implementations of generative EmCom, and investigating how language formed by embodied agents sharing similar environments could serve as an effective prior distribution for world model learning.
+In fact, recent work has already begun to demonstrate the fruitfulness of this approach; for example, Nomura et al.
+[88] have extended CPC to dynamic environments by developing a decentralized collective world model.
 From a theoretical perspective, a deeper mathematical formalization of the interaction between individual and collective learning processes will be crucial.
@@
 By pursuing these directions, we can work toward a more comprehensive understanding of language emergence and its role in environmental adaptation while advancing both the theoretical foundations and practical applications of AI systems.
+7 Conclusion
+This paper addressed the fundamental puzzle of how LLMs acquire world knowledge without embodied experience.
+We proposed the “Collective World Model” hypothesis, arguing that LLMs learn from a rich representational structure that is the product of a society-wide, interactive sense-making process.
+We provided a formal foundation for this claim with the framework of Generative EmCom, which models language emergence as a decentralized Bayesian inference process driven by the principles of CPC.
+This perspective unifies the cognitive processes of the individual with the collective evolution of language, offering a coherent explanation for the remarkable capabilities of modern AI and charting a path toward a deeper integration of communication, perception, and action in intelligent systems.
+Author Contributions
+Tadahiro Taniguchi: Conceptualization, Writing – Original Draft, Writing – Review & Editing, Supervision, Funding Acquisition.
+Ryo Ueda: Writing – Original Draft, Writing – Review & Editing, Funding Acquisition.
+Tomoaki Nakamura: Writing – Original Draft, Writing – Review & Editing, Funding Acquisition.
+Masahiro Suzuki: Writing – Original Draft, Writing – Review & Editing, Funding Acquisition.
+Akira Taniguchi: Writing – Original Draft, Writing – Review & Editing.
+Declaration of Generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the authors used Gemini 2.5 pro (Google) and ChatGPT 4o in order to improve readability and language.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
+Acknowledgement
+This work was supported by JSPS KAKENHI Grant Numbers JP21H04904, JP23H04835, JP23H04974, JP23K28181, and the JST Moonshot R&D Grant Number JPMJMS2011.